### v1

# TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning

Abstract
Language model continual learning (CL) has recently garnered significant interest due to its potential to adapt large language models (LLMs) to dynamic real-world environments without re-training. A key challenge in this field is catastrophic forgetting, where models lose previously acquired knowledge when learning new tasks. Existing methods commonly employ multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge for each task, but these approaches lack efficiency and overlook the potential for knowledge transfer through task interaction. In this paper, we present a novel CL framework for language models called Task Skill Localization and Consolidation (TaSL), which enhances knowledge transfer without relying on memory replay. TaSL first divides the model into ‘skill units’ based on parameter dependencies, enabling more granular control. It then employs a novel group-wise skill localization technique to identify the importance distribution of skill units for a new task. By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer. As a result, TaSL achieves a superior balance between retaining previous knowledge and excelling in new tasks. TaSL also shows strong generalizability, suitable for general models and customizable for PEFT methods like LoRA. Additionally, it demonstrates notable extensibility, allowing integration with memory replay to further enhance performance. Extensive experiments on two CL benchmarks, with varying model sizes (from 220M to 7B), demonstrate the effectiveness of TaSL and its variants across different settings.
Index Terms:
Language model continual learning, catastrophic forgetting, skill localization.
I Introduction
Equipping large language models (LLMs) with continual learning (CL) capabilities to sequentially learn different tasks is essential for their deployment in real-world scenarios [1, 2]. This capability enables LLMs to dynamically adapt to new tasks and acquire additional knowledge [3, 4]. An effective CL system must address two critical challenges: (1) Catastrophic Forgetting (CF)  [5], the phenomenon where a model’s proficiency in previous tasks deteriorates as it learns new ones, and (2) Knowledge Transfer (KT)  [6], enhancing task performance through the transfer of knowledge. KT can be categorized into forward transfer, which improves new task performance using knowledge from previous tasks, and backward transfer, which enhances performance on previous tasks after learning a new relevant task. Achieving a balance between retaining previous knowledge and excelling in new tasks is vital for success.
Figure 1: Conceptual illustration of TaSL. By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-specific and task-shared parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
Given the considerable computational demands, recent efforts have explored CL for LLMs using parameter-efficient fine-tuning (PEFT) methods [7, 8], such as Low-Rank Adaptation (LoRA) [7]. Traditional rehearsal-based approaches, which require storing data examples from past tasks for replay, face significant privacy and memory issues [9, 10, 3].
Another prominent line of work in recent studies is the parameter isolation CL methods [11, 12, 13, 14]. These methods assign a dedicated PEFT block for each new task to acquire task-specific knowledge, preserving the block for selective activation during testing. However, they exhibit significant limitations in effectively addressing the KT and CF challenges.
On one hand, the prevalent design of these frameworks, which utilizes multiple PEFT blocks, introduces redundancy and inefficiency [15]. The number of required PEFT blocks increases with each new task, leading to substantial memory storage costs and making it difficult to handle long task sequences [16, 17]. Moreover, such approaches essentially create separate expert models for each task, which limits their ability to generalize to unseen tasks [18] (Limitation 1).
On the other hand, these parameter-isolation CL methods do not consider interactions between tasks, which hinders KT. For instance, the method in [12] involves learning each PEFT block separately within individual tasks. Similarly, Orthogonal Low-Rank Adaptation (O-LoRA) [19] updates parameters via gradient projection in orthogonal subspaces. Although these approaches may mitigate CF to some extent, they cut off the potential transfer of knowledge stored across different PEFT blocks, thus impeding bi-directional KT among various tasks and leading to suboptimal performance (Limitation 2).
To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [20], a novel CL framework designed to improve KT between tasks without relying on memory replay. Our approach is motivated by recent findings that model parameters contribute unevenly to performance [21]. For instance, the authors in [22] uncovered a core region on LLMs crucial for all languages, while distinct monolingual regions exist for different languages. Preserving these important regions can help avoid forgetting. The research in [15] also found that during fine-tuning with LoRA, many redundant parameter modifications are often retained, suggesting that PEFT blocks contain significant redundancies.
Based on these insights, TaSL facilitates KT by identifying and consolidating the importance distribution of model parameters across tasks. TaSL initially employs a group-wise importance-aware skill localization technique that utilizes gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task. By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1. Our innovative skill consolidation phase then categorically integrates weights from previous tasks with the current one, enabling effective KT while minimizing forgetting.
In detail, TaSL first reconstructs the model or PEFT block into fine-grained “skill units”. A skill unit refers to a distinct subset of model parameters that encapsulates specific functional capabilities or knowledge relevant to a particular task, such as the Query matrix within the self-attention layer. By operating at this finer granularity, we can localize and consolidate task-specific and shared knowledge within a single PEFT block, rather than adapting a separate PEFT block for each task as in previous works (addressing Limitation 1).
The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task. Our approach, focusing on parameter space rather than dataset-driven categorization of task-specific and task-shared knowledge, offers a more effective solution to managing KT in LLMs, overcoming inaccuracies caused by dataset noise. Our skill consolidation stage, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge. During this phase, we promote forward KT by initializing new tasks with previously fine-tuned weights. For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, enhancing their capability. To prevent CF, we preserve the integrity of skill units containing previous task-specific knowledge, ensuring they remain unaffected by new task learning (addressing Limitation 2).
Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA. While the original TaSL framework has shown promising results, it has only been validated on a specific sub-task, and its effectiveness on general CL benchmarks has yet to be explored. Additionally, on the technical side, the skill localization process in TaSL relies on first-order gradients, which may not always provide precise importance localization. Furthermore, the skill consolidation phase in TaSL involves many hyperparameters, potentially complicating the averaging process.
To address these issues, we propose TasLoRA, a LoRA-tailored version of TaSL. TasLoRA reconstructs the LoRA adapter into new skill units based on parameter dependencies, ensuring more efficient management of knowledge through sequential task learning. During the fine-tuning process, an additional orthogonal loss is added to ensure that skill units within the same LoRA adapter capture distinct latent semantic features. For skill localization, TasLoRA employs a novel second-order gradient approximate group-wise metric to compute importance scores more precisely. For skill consolidation, we transition from the previous hard-mask model averaging approach to a soft-masking strategy, using an adaptive averaging technique. This allows for a more flexible integration of task-specific and shared parameters, dynamically adjusting the contribution of each skill unit based on its importance.
Furthermore, to adapt to a broader range of application scenarios, we combine TaSL with memory replay, introducing the TaSL-M model. Through extensive experiments on two CL benchmarks with different parameter-level backbones (from 220M to 7B), our TaSL framework, along with its variants, excels in mitigating CF and showcases remarkable capabilities for KT, outperforming state-of-the-art (SOTA) methods.
The main contributions are summarized as follows:
  • We propose a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning. By identifying and consolidating task-specific and task-shared knowledge at a granular skill unit level, TaSL achieves effective knowledge transfer and mitigates catastrophic forgetting, overcoming the limitations of previous approaches.
  • We develop various group-wise skill localization and fine-grained skill consolidation techniques. For instance, the parameter importance metric in skill localization can be based on first-order gradients or a new second-order gradient approximation method. In skill consolidation, our model averaging strategies include a categorical hard-mask approach and an adaptive soft-mask method.
  • The TaSL framework demonstrates strong generalizability and extensibility. The flexible design of skill units enables TaSL to be easily tailored to PEFT methods, such as LoRA, optimizing performance for specific model architectures. Additionally, TaSL can be integrated with memory replay technique to further enhance performance and adapt to a broader range of application scenarios.
  • Extensive evaluation on two CL benchmarks underscores the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios. Furthermore, TaSL consistently excels across diverse model sizes (ranging from 220M to 7B), different model architectures (T5 and LLaMA-2), and unseen tasks.
II Related Work
II-A Continual Learning
Continual learning [3] aims to develop learning algorithms capable of accumulating knowledge from non-stationary data.
Conventional Continual Learning methods can be divided into three categories: (i) Regularization-based methods add explicit regularization terms to preserve the knowledge of previous tasks [13, 23]. (ii) Rehearsal-based methods address catastrophic forgetting by storing old training samples [24, 25] or by training generative models to provide pseudo samples of previous tasks [26, 27]. Due to the memory and data privacy issues associated with these methods, we focus on rehearsal-free CL methods. (iii) Parameter isolation-based methods dynamically expand model capacity or assign isolated parameters dedicated to each task throughout the CL process to prevent interference between tasks [28, 29, 30].
Language Model Continual Learning with PEFT. Based on PEFT methods, current approaches for the continual learning of LLMs adopt the concept of parameter isolation, using a pipeline fashion to learn and select PEFT blocks for each task [12, 31, 32]. However, these approaches are limited in effectively addressing the challenges of catastrophic forgetting (CF) and knowledge transfer (KT) [33]. For instance, the authors in [19] and [34] constrain the learning of PEFT blocks to maintain orthogonality, restricting KT among different tasks. Although the recent SAPT method [16] and DAP method [35] achieves KT, it relies on pseudo sample replay or unlabeled domain corpora.
Our method achieves KT without memory replay, offering unique advantages in data privacy and parameter efficiency. By using layering low-rank adapters on the key and value projection matrices of transformer blocks, we balance retaining previous knowledge and excelling in new tasks.
II-B Skill Localization
Research shows that model parameters do not all contribute equally to performance [36]. The authors in [21] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, suggesting that fine-tuning only these critical parameters can achieve results comparable to full model fine-tuning. However, their method involves additional steps to identify and retrain these key parameters during the post-fine-tuning phase, which reduces efficiency.
Inspired by techniques from the pruning community, previous studies have utilized gradient-based metrics to pinpoint important parameters during fine-tuning. Sensitivity-based scoring [37, 38] evaluates the impact on training loss, and sensitivity smoothing, as employed by [39], helps remove unnecessary parameters for more efficient fine-tuning. Nonetheless, these methods often result in element-wise pruning, where individual parameters are removed based on their importance scores. This approach can be computationally and storage-intensive, as it requires detailed management of numerous parameters throughout the model.
Building on these advancements, we propose a novel importance-aware skill localization method. This approach uniquely differentiates between task-specific and shared parameters to mitigate forgetting and achieve KT in CL.
III Preliminaries
III-A Continual Learning
Continual learning [40] aims to create learning algorithms that can continuously accumulate knowledge from ongoing sequences. In supervised continual learning, a sequence of tasks $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{K}\}$ is presented in a streaming manner. Each task $\mathcal{T}_{k}$ includes a distinct target dataset $\mathcal{D}_{k}=\left\{\left(x_{i}^{k},y_{i}^{k}\right)\right\}_{i=1}^{N_{k}}$ of size $N_{k}$, where $x_{i}^{k}\in\mathcal{X}_{k}$ and $y_{i}^{k}\in\mathcal{Y}_{k}$. The model must adapt to these tasks sequentially, having access only to $\mathcal{D}_{k}$ during the $k$-th task. Generally, with a prediction model $f$ parameterized by $\Theta$, continual learning aims to optimize the following objective across all tasks:
$\max_{\Theta}\sum_{k=1}^{K}\sum_{x,y\in\mathcal{D}_{k}}\log p_{\Theta}(y\mid x)$ (1)
The notation $f_{k}$ refers to the model after training on task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ denotes the model after averaging for $\hat{f}_{k-1}$ and $f_{k}$. Our TaSL framework aims to address a more challenging scenario where the model cannot access any historical data during training [19]. Nevertheless, to adapt to a wider range of application scenarios, we have also combined TaSL with memory replay. In this setting, we randomly save $\left|\mathcal{M}\right|$ samples from the training set of each previous task $\mathcal{T}_{i}$ in memory $\mathcal{M}_{i}$ and jointly train the model on new task data $\mathcal{D}_{k}$ and memory $\mathcal{M}_{<k}$. This extension, called TaSL-M, leverages memory replay to enhance performance.
III-B Low-Rank Adaption
Low-Rank Adaption (LoRA) [7] is a parameter-efficient fine-tuning method to adapt LLMs to novel tasks. It assumes that the changes of parameters lie in a low-rank space when the model is fully fine-tuned on a downstream task. This approach eliminates the necessity of fine-tuning the entire model, thereby improving computational efficiency and resource utilization.
Inspired by the low-rank internal dimensionality [41], LoRA hypothesizes the updates to the weights also has a low “intrinsic rank” during adaptation. For a pre-trained weight matrix $W^{(0)}\in\mathbb{R}^{out\times in}$ that takes $x$ as input, LoRA modifies the output $h$ of $W^{(0)}$ with a low-rank decomposition:
$h=W^{(0)}x+\Delta Wx=W^{(0)}x+BAx,$ (2)
while $B\in\mathbb{R}^{out\times r}$, $A\in\mathbb{R}^{r\times in}$, and the rank $r\ll\min(in,out)$. $A$ is initialized from a random Gaussian distribution and $B$ is initialized with all zeros. During the training stage, $W^{(0)}$ is fixed, and only $BA$ is updated. Due to its superior performance, LoRA has become one of the most popular PEFT methods in NLP community.
IV Proposed Method: TaSL
IV-A Overview
TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to accurately identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs a novel fine-grained model averaging strategy to integrate model weights from both current and past tasks for effective knowledge transfer.
Figure 2 provides a comprehensive overview of our proposed TaSL framework, with the following subsections detailing each component and the corresponding extensions.
Figure 2: Left: Illustration of reconstructing LoRA as fine-grained skill units. Right: Overview of TaSL. Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware skill localization method during fine-tuning. Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$. The integration is guided by the importance distributions of skill units across various tasks. This process can be repeated with the introduction of each subsequent task.
IV-B Fine-Grained Skill Unit
Before we localize the importance of parameters, we need to define a structure that helps us better organize the knowledge stored in the model. This structure will provide a clearer understanding of how to mitigate CF and achieve KT. Therefore, we refer to the basic unit of stored skills or knowledge in the model as a “skill unit.” Based on different usage scenarios, we offer the following two structures for skill units:
IV-B1 Matrix-Level Skill Unit
In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA. This approach aligns with the original TaSL framework. The advantage of this division is that it is model-agnostic, making it applicable to both traditional full-parameter fine-tuning and the latest parameter-efficient fine-tuning methods, such as LoRA. This provides a higher degree of generalizability.
IV-B2 LoRA-Tailored Skill Unit
While the matrix-level division is simple and intuitive, it did not sufficiently consider the intra-matrix redundancy and inter-matrix dependencies. To address these limitations and the inefficiencies noted in recent work [19], which treats each LoRA adapter as a container for task-specific knowledge, we decompose the model matrices into new, finer-grained “skill units” based on parameter dependencies tailored for LoRA. The construction process for these LoRA-tailored skill units is as follows.
As shown in Eq (2), $A$ and $B$ can be viewed as a combination of a set of vectors: $A=[a_{1},a_{2},\cdots,a_{r}]$, $B=[b_{1},b_{2},\cdots,b_{r}]$, where $a_{i}\in\mathbb{R}^{in}$, $b_{i}\in\mathbb{R}^{out}$. Thus $BA$ can be further disassembled as:
$W$ $=W^{(0)}+b_{1}a_{1}+b_{2}a_{2}+\cdots+b_{r}a_{r}$ (3)
$=W^{(0)}+u_{1}+u_{2}+\cdots+u_{r},$
where skill unit $u_{i}$ is a matrix obtained by the product of the vectors $b_{i},a_{i}$. Therefore, LoRA can be viewed as a fusion of knowledge learned from multiple skill units:
$W$ $=W^{(0)}+\sum_{i=1}^{r}b_{i}a_{i}=W^{(0)}+\sum_{i=1}^{r}u_{i}$ (4)
This division strategy effectively decomposes the parameters of a single LoRA adapter, reducing the redundancy in the traditional approach of employing separate LoRA adapters for each task [19]. To ensure that different skill units within the same layer learn distinct latent semantic features, we apply a regularization term as introduced by [39]:
$R(A,B)=\|A^{T}A-I\|_{F}^{2}+\|B^{T}B-I\|_{F}^{2},$ (5)
where $I$ is an identity matrix, this forces $A$ and $B$ to be orthogonal after training.
IV-C Importance-aware Skill Localization
To calculate the importance of each skill unit $u$, we introduce a group-wise metric that also addresses the significant computational and storage burdens associated with previous parameter-level importance calculation methods [42]:
$\mathcal{I}(u)=\frac{1}{in\times out}\sum\limits_{i=1}^{in}\sum\limits_{j=1}^{out}s(w_{ij})$ (6)
where $w_{ij}$ denotes the trainable parameters, and $in\times out$ represents the total parameter count in a skill unit $u$. $\mathcal{I}(u)$ measures the collective importance of all parameters within each skill unit, where higher values signify increased importance. The importance function $s(\cdot)$ for individual parameters, inspired by the pruning community [43], is quantified by measuring the impact of its removal on the loss. Specifically, to estimate the importance of $W_{i}$, we calculate the change in loss when the parameter is zeroed out using:
$I_{W_{i}}=|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}_{W_{i}}(\mathcal{D})-\mathcal{L}_{W_{i}=0}(\mathcal{D})|$ (7)
$=\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W_{i}}W_{i}-\frac{1}{2}W_{i}H_{ii}W_{i}+\mathcal{O}(\|W_{i}\|^{3})\right|$
where $H$ is the hessian matrix, $W_{i}$ represents the $i$-th parameter in $W$, and $\mathcal{L}$ represents the next-token prediction loss. If removing a parameter has a significant influence, then the model is sensitive to it, and we should retain it [44]. Based on Eq. (7), we can derive the following two metrics for calculating parameter importance:
IV-C1 First-Order Gradient-Based Metric
In original TaSL framework, we followed the standard approach, defining $s(\cdot)$ as the magnitude of the gradient-weight product:
$I_{W_{i}}=\left|W_{i}\nabla_{W_{i}}\mathcal{L}\right|$ (8)
This metric only considers the first-order gradient term in the importance calculation. However, the importance of a parameter is also reflected in the second-order gradients (Hessian matrix). Ignoring second-order gradient information may lead to biases in localization, thus affecting model performance. Therefore, we propose the following new metric.
IV-C2 Second-Order Gradient Approximation Metric
Direct computation of the Hessian matrix for LLMs is impractical due to its $\mathcal{O}(N^{2})$ complexity. To address this, we propose a second-order gradient approximation for the importance metric, which balances the efficiency of first-order methods with the precision of second-order gradients. To reduce the computational complexity, the diagonal of the Hessian $H_{ii}$ can be approximated by the Fisher information matrix [45], and the importance can be defined as:
$I_{W_{i}}\approx\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W_{i}}W_{i}-\frac{1}{2}\sum_{j=1}^{N}\left(\frac{\partial\mathcal{L}(\mathcal{D}_{j})}{\partial W_{i}}W_{i}\right)^{2}\right|$ (9)
However, calculating the importance as specified in Eq. (9) across the entire training set introduces significant challenges, as model training typically only has access to mini-batch data. This means the metric is subject to variability due to stochastic sampling and training dynamics, introducing large uncertainty in estimating sensitivity. To mitigate this, we propose a more reliably importance metric based on sensitivity smoothing and uncertainty quantification [39]:
$\begin{split}\bar{I}^{(t)}\left(w_{ij}\right)=\alpha_{1}\bar{I}^{(t-1)}\left(w_{ij}\right)+\left(1-\alpha_{1}\right)I^{(t)}\left(w_{ij}\right)\end{split}$ (10)
$\begin{split}\bar{U}^{(t)}\left(w_{ij}\right)=\alpha_{2}\bar{U}^{(t-1)}\left(w_{ij}\right)+\\ \left(1-\alpha_{2}\right)\left|I^{(t)}\left(w_{ij}\right)-\bar{I}^{(t)}\left(w_{ij}\right)\right|\end{split}$ (11)
where $\alpha_{1}$ and $\alpha_{2}$ are smoothing factors, and $t$ is the iteration number. $\bar{I}^{(t)}$ represents smoothed sensitivity and $\bar{U}^{(t)}$ is the uncertainty term quantified by the local variation between $I^{(t)}$ and $\bar{I}^{(t)}$. Using the exponential moving average of the importance metric, we can retain and explore the trajectory gradient for a longer time, providing a more robust and precise importance assessment. Importance is then determined by:
$s^{(t)}\left(w_{ij}\right)=\bar{I}^{(t)}\left(w_{ij}\right)\cdot\bar{U}^{(t)}\left(w_{ij}\right)$ (12)
To compute the importance score of each skill unit for current task $\mathcal{T}_{k}$, we employ Eq. (6) during fine-tuning. The model $f$ with $n$ skill units is denoted as $\mathcal{U}=\left\{u_{1},\ldots,u_{n}\right\}$, with their importance scores for task $\mathcal{T}_{k}$ denoted by $\mathcal{I}(\mathcal{U}_{k})\in\mathbb{R}^{n}$. The detailed computation process is provided in Algorithm 1.
Algorithm 1 Importance-aware Skill Localization
0:  Training dataset $\mathcal{D}_{k}$ for task $\mathcal{T}_{k}$; total training iterations $T$; hyperparameters $\alpha_{1},\alpha_{2}$.
  for $t=1,\ldots,T$ do
     Sample a mini-batch from $\mathcal{D}_{k}$ and compute the gradient $\nabla\mathcal{L}$;
     Compute the sensitivity $I\left(w_{ij}\right)$ via Eq. (8) or Eq. (9);
     Update $\bar{I}^{(t)}$ via Eq. (10) and $\bar{U}^{(t)}$ via Eq. (11);
  end for
  Compute the importance score $\mathcal{I}(u^{k}_{i})$ for each skill unit $u^{k}_{i}$ by Eq. (6), for $i=1,\ldots,n$.
  $f_{k}$ and importance scores $\mathcal{I}(\mathcal{U}_{k})$ for $\mathcal{U}_{k}$.
After computing importance scores for each skill unit at the current task $\mathcal{T}_{k}$, it is essential to compare these with scores from all previously learned tasks to distinguish between task-specific and task-shared skill units. To avoid the inefficiency of storing scores for each past task, we aggregate importance scores from all prior tasks into a cumulative score for tasks up to $\mathcal{T}_{k-1}$. This method allows for the iterative refinement of accumulated scores without separately saving past task scores. The skill units with these cumulative scores up to $\mathcal{T}_{k-1}$ are denoted as $\hat{\mathcal{U}}_{k-1}$, calculated using:
$\begin{split}\mathcal{I}(\hat{\mathcal{U}}_{k-1})=\beta\text{Norm}(\mathcal{I}(\hat{\mathcal{U}}_{k-2}))+(1-\beta)\text{Norm}(\mathcal{I}(\mathcal{U}_{k-1}))\end{split}$ (13)
where $\beta\in[0,1]$, and Norm($\cdot$) normalizes importance scores to the [0, 1] range, thus resolving discrepancies across models. The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set to be equal to $\mathcal{I}(\mathcal{U}_{1})$. Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
IV-D Skill Consolidation
After skill localization, the subsequent vital phase involves consolidating the knowledge in the skill unit into a unified framework. This process demands a sophisticated model averaging approach considering various factors to optimize task performance. Traditional coarse-grained model averaging assumes that all model weights are equally important for the training task [46, 47], which can be written as the following iterative computation format:
$\hat{f}_{k}=\lambda\hat{f}_{k-1}+\left(1-\lambda\right)f_{k}$ (14)
However, coarse-grained methods may overemphasize weights that are irrelevant to the current task, contaminating previously acquired task-specific knowledge and leading to forgetting. To address this issue, we introduce two fine-grained averaging strategies that focus on skill units rather than the entire model. Our approach distinguishes between task-shared and task-specific skill units, applying weighted averaging to parameters within each unit. This tailored method provides a customized solution to meet the distinct needs of each task.
IV-D1 Static Weighted Consolidation
This strategy, used in the original TaSL framework, begins by setting importance thresholds $\delta$ using quantiles to select the top 20% of skill units based on importance scores. A skill unit $u_{i}^{k}$ is deemed important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ is above $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
Our static weighted consolidation strategy customizes parameter combination for each skill unit, based on its importance under different tasks, as follows:
$\hat{u}_{i}^{k}=\begin{cases}\gamma\hat{u}_{i}^{k-1}+(1-\gamma)u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{+}$}\\ \hat{u}_{i}^{k-1},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{-}$}\\ u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{+}$}\\ \frac{1}{2}(\hat{u}_{i}^{k-1}+u_{i}^{k}),&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{-}$}\\ \end{cases}$ (15)
This approach performs the element-wise adjustment of parameters within each skill unit based on its relevance to previous and current tasks, using hyperparameter $\gamma$ to control their influences.
IV-D2 Adaptive Weighted Consolidation
Building upon the original TaSL framework, we propose an Adaptive Weighted Consolidation strategy as our latest extension. In contrast to the manual setting of multiple hyperparameters required in the original method, this adaptive strategy simplifies the process by automatically adjusting to various scenarios, making the averaging process more efficient and flexible. In detail, for a specific skill unit $u_{i}$, the weighting coefficients are:
$\begin{split}\hat{\lambda}^{k-1}_{i}=\exp\left(\mathcal{I}(\hat{u}^{k-1}_{i})/\tau\right),\lambda^{k}_{i}=\exp\left(\mathcal{I}(u^{k}_{i})/\tau\right)\end{split}$ (16)
where both $exp$ and temperature coefficient $\tau$ are scaled to the raw importance score. Then the updated model parameters are:
$\begin{split}\hat{u}_{i}^{k}=\left(\frac{\hat{\lambda}^{k-1}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot\hat{u}_{i}^{k-1}+\left(\frac{\lambda^{k}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot u_{i}^{k}\end{split}$ (17)
By using Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high while $\mathcal{I}(u^{k}_{i})$ is low, it indicates that the skill unit $u_{i}$ is more important for the history task. According to Eq. (15) or (17), we maintain the previous task-specific knowledge stored in $\hat{u}^{k-1}_{i}$ untouched to prevent the contamination of historical knowledge with task-irrelevant information in $u^{k}_{i}$, thereby mitigating CF.
  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is low and $\mathcal{I}(u^{k}_{i})$ is high, it suggests that the skill unit $u_{i}$ is more important for the current task. Since the model’s training initializes from the endpoint of the previous task, the historically learned knowledge is utilized to enhance performance on the current task. So we ensure the integrity of the parameters in $u^{k}_{i}$, which stores current task-specific knowledge, thereby facilitating forward KT.
  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are high, it indicates that the skill unit $u_{i}$ is important for both the previous and current tasks. Consequently, we integrate newly acquired knowledge into this task-shared skill unit to enable backward KT.
  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are low, it indicates the skill unit is not significantly relevant to any task. In this case, simply averaging the parameters within this unit is sufficient.
Skill consolidation is performed before starting a new task in CL, utilizing the averaged model for subsequent task initialization. Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are retained for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$. Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
Algorithm 2 TaSL
0:  Dataset $\mathcal{D}_{k}$ for task $k$ = $1,\ldots,K$; initial pre-trained model $f_{0}$; hyperparameters $\alpha_{1},\alpha_{2},\beta,\tau$.
1:  # sequential tasks.
2:  for task $k$ = $1,\ldots,K$ do
3:     # skill localization.
4:     Get $f_{k}$ and calculate $\mathcal{U}_{k}$ by Algorithm (1);
5:     if $k=1$ then
6:        # initialization at beginning task.
7:        $\hat{f}_{1}\leftarrow f_{1}$, $\hat{\mathcal{U}}_{1}\leftarrow\mathcal{U}_{1}$;
8:     else
9:        # skill consolidation.
10:        for skill unit $i$ = $1,\ldots,n$ do
11:           Calculate $\hat{u}_{i}^{k}$ by Eq. (17);
12:        end for
13:        Get the averaged model $\hat{f}_{k}$ based on $\hat{\mathcal{U}}_{k}$;
14:        Calculate accumulated importance score $\mathcal{I}(\hat{\mathcal{U}}_{k})$ according to Eq. (13);
15:     end if
16:  end for
V TaSL with Memory Replay: TaSL-M
To adapt TaSL to scenarios where historical data can be used, we propose a memory replay-enhanced version, TaSL-M, to further improve model performance. The implementation process is shown in Figure 3.
Figure 3: Overview of TaSL-M.
After training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$. Before the next task arrives, we introduce a replay stage, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$. This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge. Specifically, during the replay stage, we do not fine-tune the parameters of all skill units using historical data, as this could negatively impact the performance on the current task. Instead, we fix the current task-specific skill units and fine-tune only the remaining parameters. This approach achieves a better balance between retaining previous knowledge and excelling in new tasks.
VI Experiments and Analysis
VI-A Experimental Setup
VI-A1 Datasets
We utilize the SuperNI Benchmark [48], a comprehensive benchmark designed to evaluate diverse NLP tasks using expert-written instructions. This benchmark facilitates thorough evaluation in more practical settings for the CL of LLMs. It includes tasks in dialogue generation, information extraction, question answering, summarization, and sentiment analysis. Following the established CL setup [16], three tasks are chosen from each type, creating a sequence of 15 tasks in total for evaluation. For each task, 1,000 instances from the dataset are randomly selected for training, and 100 instances are used for validation and testing.
Additionally, we employ the Long Sequence Benchmark [18], a continual learning benchmark consisting of 15 classification datasets. In line with [19], we randomly select 1,000 samples for training each task and reserve 500 samples per class for validation and testing. We explore two different task orders for each benchmark. Please refer to the Appendix for more details about the tasks and orders.
TABLE I: The overall results on two continual learning benchmarks with T5-large model. All results are averaged over two different orders of each benchmark. Our rehearsal-free TaSL outperforms the previous best method, O-LoRA, by achieving an average 2.3% absolute improvement on AP, an average 2.8% absolute improvement on FWT, and an 4.8% absolute increase in BWT.
Method Memory-Replay SuperNI Benchmark Long Sequence Benchmark
[…]
SeqLoRA 6.43 1.58 -30.94 9.72 6.81 -73.37
IncLoRA 19.12 2.03 - 31.24 62.50 2.62 -15.34
ProgPrompt [18] 3.34 5.29 -33.18 7.98 6.63 -66.71
EWC [46] 17.46 4.20 -28.61 45.45 3.73 -25.93
L2P [49] 12.73 1.14 -7.95 57.98 8.36 -16.63
LFPT5 [50] 24.76 7.46 -24.41 67.01 9.48 -12.80
O-LoRA [19] 25.89 8.14 -24.59 69.24 10.15 -4.05
TaSL (ours) 26.41 11.78 -18.55 70.71 11.80 -3.27
TasLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -2.04
Replay 35.37 2.35 -15.79 75.28 3.28 -1.88
SAPT [16] 51.54 8.88 -0.57 82.02 9.86 -1.25
TaSL-M (ours) 52.12 11.13 -1.31 84.26 12.32 -1.08
TasLoRA-M (ours) 53.01 11.21 -0.81 84.33 11.71 -0.98
VI-A2 Evaluation Metrics
We denote $a_{j,i}$ as the testing performance (Accuracy for classification task and Rouge-L [51] for others) on the $i$-th test set of task right after training on $j$-th task. The performance of CL is assessed using three metrics:
$\mathbf{AP}=\frac{1}{K}\sum_{i=1}^{K}a_{K,i}$ (18)
$\mathbf{FWT}=\frac{1}{K-1}\sum\limits_{i=2}^{K}a_{i-1,i}$ (19)
$\mathbf{BWT}=\frac{1}{K-1}\sum\limits_{i=1}^{K-1}a_{K,i}-a_{i,i}$ (20)
Average Performance (AP) [52] calculates the average performance across all tasks after training on the final task. Forward Transfer (FWT) [53] evaluates a model’s generalization ability by measuring the averaged zero-shot performance. Backward Transfer (BWT) [54] assesses the impact of new learning on previous tasks. Negative BWT indicates the model lost some previously acquired knowledge.
VI-A3 Baselines
We evaluate TaSL against the following PEFT-based continual learning baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders. IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks. Replay: replays real samples from old tasks when learning new tasks to avoid forgetting. EWC [46]: finetune the model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks. L2P [49]: uses the input to dynamically select and update prompts from a fixed prompt pool. LFPT5 [50]: continuously trains a soft prompt for each task with generative replay and an auxiliary loss. Prog-Prompt [18]: sequentially concatenates previous learned prompts to the current one during the training and testing time. O-LoRA [19]: learns tasks in different LoRA subspaces that are kept orthogonal to each other and sums all LoRA weights up at testing time. SAPT [16]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
Table II compares our TaSL with these baselines, highlighting three distinct advantages: data privacy-friendliness, model parameter-friendliness, and generalization-friendliness. To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our base framework utilizes original components, including matrix-level skill units, first-order gradient-based metric for skill localization (Eq. 8), and static weighted consolidation (Eq. 15). TasLoRA: as the latest extension of TaSL, it employs LoRA-tailored skill units, the second-order gradient approximation metric for skill localization (Eq. 9), and adaptive weighted consolidation (Eq. 17). TasL-M and TasLoRA-M: these variants combine TaSL or TasLoRA with memory replay techniques to further enhance performance.
TABLE II: The comparison between TaSL and other CL methods. Specifically, RF indicates whether the method is rehearsal-free. PE indicates whether the method is parameter efficient. UT indicates whether the method can be applied to solve unseen tasks. KT indicates whether the method enables knowledge transfer.
Method RF PE UT KT
EWC [46]
OGD [55]
LFPT5 [50]
L2P [49]
EIP [12]
O-LoRA [19]
MoCL [56]
SAPT [16]
TaSL
VI-A4 Implementation Details
We utilize two language models adopted by the previous lines of works in CL for NLP: encoder-decoder T5 model [57] and decoder-only LLaMA model [58]. In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq. (10) and Eq. (11) are set to 0.85. We set $\beta$ in Eq. (13) to 0.7 and $\tau$ in Eq. (16) to 0.15. Following [59], the volume of replay samples is set to 2% of the original training set for all replay-based baseline methods. For different backbones, we utilized the following hyperparameters:
  • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, rank = 8, targeting modules [q, v] and 10 epochs.
  • LLaMA (7B): With a learning rate of 3e-4, batch size of 128, a cutoff length of 512, and 10 epochs. Lora settings were rank = 8, alpha = 16, dropout = 0.05, targeting modules [q_proj, v_proj]. For testing, settings included temperature = 0.02, top_p = 0, top_k = 1, max new tokens = 128.
Experiments are carried out using 4 NVIDIA A100.
VI-B Main Results
Overall CL results of different methods at the same T5-large backbone are summarized in Table I.
Firstly, TaSL and its variants demonstrate superior CL performance through effective knowledge transfer. Compared to previous rehearsal-free methods such as IncLoRA and O-LoRA, TaSL excels in addressing the two critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
Moreover, our enhanced version, TasLoRA, shows significant improvements over TaSL. Specifically, TasLoRA achieves average gains of 1.3% on AP and 1.4% on BWT, indicating that our extensions to each component are effective. Notably, even without relying on memory replay, our method nearly matches the performance of SAPT with memory replay on Long Sequence Benchmark, particularly in BWT, with only a slight difference of 0.8% (-2.04% vs. -1.25%).
For memory-based methods, both TaSL-M and TasLoRA outperform the best SAPT methods, further demonstrating the versatility and robustness of our framework. For the convenience of subsequent experimental analysis and comparison, we focus on the performance of TasLoRA, the best-performing memory-free version, for further evaluations.
Secondly, TasLoRA consistently demonstrates superior performance across various backbones. To further validate the effectiveness of our framework, we conducted experiments using a range of parameter-level backbones, as shown in Figure 4, which highlight performance improvements with increasing model size. Across different backbones, our method consistently outperforms traditional approaches. For example, in T5-xl, TasLoRA significantly increases the AP metric from 28.6% to 33.3%, and achieves substantial improvements in both FWT and BWT metrics, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively. These results further validate the robust generalization ability of our proposed framework.
(a) Results on the SuperNI benchmark
(b) Results on the Long Sequence benchmark
Figure 4: Performance of TaSL w/ different backbones.
Thirdly, adaptive skill consolidation effectively mitigates catastrophic forgetting. To rigorously assess our model’s effectiveness in countering forgetting, we analyzed its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark. Figure 5 shows that our TasLoRA method significantly slows the forgetting rate, with an average performance decrease of only 11% after training on the final task. In stark contrast, vanilla backbones exhibit a substantial average performance drop of 20%, underscoring our method’s superior ability to mitigate forgetting.
Figure 5: Performance trajectory of Task 1 on Long Sequence Benchmark during the CL process.
VI-C Performance on Unseen Tasks
Following previous work [16], we further select three tasks from each one of the above task categories to assess the TaSL’s cross-task generalization ability. This is also a crucial dimension for evaluating CL algorithms. Table III shows the results. T5-ZS represents the zero-shot approaches for task adaptation, respectively. TasLoRA achieved the best performance, which can be attributed to its ability to identify task-specific and shared parameters effectively.
TABLE III: Results on unseen tasks based on the T5-Large backbone model. We report the average Rouge-L of the 3 tasks under each category.
Unseen Tasks Avg.
Dialog IE QA Sum SA
T5-ZS 7.49 6.70 4.28 12.14 4.54 7.03
O-LoRA 4.39 9.89 25.38 8.26 50.41 19.67
LFPT5 6.96 35.32 35.00 13.26 21.51 22.41
TasLoRA 10.32 31.34 37.13 14.20 47.17 28.03
VI-D Visualization of Skill Units
We visualized the distribution of importance scores for the skill units across tasks on T5-large, as shown in Figure 6 and Figure 7, leading to several critical insights:
  • There is a noticeable variation in the importance of skill units for the same task, with important skill units making up only a small fraction of all trainable LoRA parameters.
  • The distribution of important skill units is task-dependent, indicating both task-shared and specific parameters, confirming TaSL’s validity.
  • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more significant role. In contrast, for generative tasks, such as dialogue generation and summarization in the SuperNI benchmark (Figure 7), both the encoder and decoder skill units are important, with both the lower and upper layers of the network being crucial.
  • Within each layer, the importance of the Query (Q) matrices in the attention mechanism consistently surpasses that of the Value (V) matrices.
Figure 6: Visualization of importance scores for skill units across different tasks on T5-large for the Long Sequence Benchmark.
Figure 7: Visualization of importance scores for skill units across different tasks on T5-large for the SuperNI Benchmark.
VI-E Ablation Study
In this section, we evaluate the impact of importance-aware skill localization, fine-grained skill consolidation, and hyperparameter sensitivity, with the results discussed below.
VI-E1 Effect of the proposed importance metric in skill localization
Our method calculates importance scores using Eq. (6). As shown in Table IV, we explore alternative importance scoring approaches: (i) to test the effectiveness of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq. (6) to include only sensitivity, as described in Eq. (9); and (ii) to assess the validity of our proposed approximate second-order gradient importance metric in Eq. (6), we use a first-order Taylor expansion to approximate the importance $I$ by using Eq. (8). The results indicate that using moving averages for importance scoring outperforms the alternatives, with the other two variants leading to performance decreases of up to 3.3%, 2.5%, and 4.2% across the three metrics. This underscores the value of accurate skill localization in enhancing model performance.
TABLE IV: Ablation study. Evaluating the impact of importance scoring variations on skill localization.
Method AP FWT BWT
vanilla T5-large 54.10 3.32 -29.63
$s\left(\cdot\right)=I\left(\cdot\right)$ 70.48 10.39 -3.81
$s\left(\cdot\right)=\left|\nabla_{w_{ij}}\mathcal{L}\right|$ 68.82 10.80 -6.22
TasLoRA (ours) 72.29 12.89 -2.04
VI-E2 Effect of Adaptive Skill Consolidation
We compared our fine-grained skill unit-level adaptive LoRA averaging mechanism against two coarse-grained strategies: (i) Weight-Ensemble, which uses a global weight to uniformly average the LoRA parameters, i.e., Eq. (14), and (ii) Exponential Moving Average (EMA) [60], which applies a running average of parameters at each fine-tuning iteration.
For Table V, Weight-Ensemble significantly improves upon the vanilla model, highlighting the benefits of coarse-grained averaging for continual learning. EMA generally outperforms Weight-Ensemble but falls short of our fine-grained approach due to its overuse of averaging, with frequent parameter adjustments within the same task potentially leading to suboptimal outcomes. Our method, which averages weights only after each task, enhances computational efficiency.
TABLE V: Ablation study. Comparing coarse- and fine-grained model averaging methods on skill consolidation.
Method AP FWT BWT
vanilla T5-large 54.10 3.32 -29.63
Weight-Ens. 63.28 7.71 -11.82
EMA 62.76 8.23 -9.80
TasLoRA (ours) 72.29 12.89 -2.04
VI-E3 Sensitivity Analysis for Hyperparameters
The proposed framework incorporates three key hyperparameters, including the $\alpha$ for computing importance scores in Eq. (10) and Eq. (11), the $\beta$ for calculating cumulative importance scores in Eq. (13), and the $\tau$ for performing weighted averaging within skill units as outlined in Eq. (16). Our analysis aims to assess the impact of varying these hyperparameters on our method’s performance, testing on the T5-large backbone model.
As evidenced in Table VI, we determine that the optimal setting for $\alpha$ is 0.55. An $\alpha$ value too low results in a performance decline, indicating that the calculated importance scores are not sufficiently accurate. Furthermore, as depicted in Table VII, we also find that $\beta$ values within a normal range do not significantly affect performance. However, excessively high or low values for $\beta$ may skew the model towards favoring either past or current task knowledge, thereby disrupting the desired balance. Nonetheless, the model’s performance remains relatively stable across most conditions, indicating a low sensitivity to hyperparameter variations.
The VIII below shows the model’s performance under different temperature coefficients, $\tau$. It can be seen that setting a higher $\tau$ (higher than 0.5) smooths the weights, thereby increasing the weight of unimportant skill units, which reduces the model’s effectiveness and may contaminate historical knowledge, leading to forgetting. On the other hand, setting a lower $\tau$ (less than 0.1) may cause larger differences in weight coefficients, thereby limiting KT between different tasks.
TABLE VI: Performance comparisons of TaSL equipped with different $\alpha$ on Long Sequence Benchmark at task order 1.
$\alpha_{1}$,$\alpha_{2}$ AP FWT BWT
Vanilla T5-large 51.6 5.6 -29.7
0.15 71.8 13.7 -3.7
0.35 71.2 14.1 -4.3
0.55 72.8 12.6 -2.5
0.85 70.7 12.9 -3.3
0.95 71.7 14.0 -3.6
TABLE VII: Performance comparisons of TaSL equipped with different $\beta$ on Long Sequence Benchmark at task order 1.
$\beta$ AP FWT BWT
Vanilla T5-large 51.6 5.6 -29.7
0.1 73.8 13.6 -3.7
0.3 73.5 12.4 -3.4
0.5 74.3 13.5 -2.2
0.7 72.7 12.9 -2.3
0.9 60.2 14.2 -5.0
TABLE VIII: Performance comparisons of TaSL (using T5-large as the backbone) equipped with different $\tau$ on Long Sequence Benchmark at task order 1.
temperature coefficient $\tau$ AP FWT BWT
0.05 69.0 14.3 -3.4
0.15 73.4 12.8 -2.1
0.25 72.2 13.2 -2.5
0.35 72.1 13.6 -2.1
0.45 72.7 12.5 -3.0
0.55 70.9 11.6 -4.2
VII Conclusion
In this paper, we introduce a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning, facilitating effective knowledge transfer across tasks. Our framework leverages an importance-aware skill localization and a fine-grained skill consolidation technique to differentiate between task-specific and shared knowledge within each skill unit, thereby mitigating forgetting. TaSL also demonstrates strong generalizability and extensibility, with optimizations to various components and the ability to integrate with memory replay methods for further performance enhancement. Comprehensive experiments showcase our framework’s exceptional ability to balance preserving past knowledge and excelling in new tasks, surpassing previous state-of-the-art methods.
While our method significantly improves efficiency and performance in large language model continual learning, some limitations remain. For instance, the choice of importance thresholds after determining the importance distribution can affect the effectiveness of subsequent consolidation. Dynamically selecting thresholds based on data distribution could lead to more accurate classification of task-shared and task-specific parameters, thereby enhancing performance. Additionally, we could explore merging the skill localization and consolidation stages, allowing for parameter consolidation based on importance during model training. This would enable more flexible adaptation and address scenarios in online continual learning.
[Dataset Details] Table IX & X show details of the datasets we used for our experiments, along with their evaluation metrics. Overall, in SuperNI, we choose 3 tasks from dialogue generation (Dialog), information extraction (IE), question answering (QA), summarization (Sum) and sentiment analysis (SA), respectively.
For the Long Sequence benchmark, this includes five tasks from the standard CL benchmark (AG News, Amazon reviews, Yelp reviews, DBpedia and Yahoo Answers), four from GLUE benchmark (MNLI, QQP, RTE, SST2), five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB movie reviews dataset. We report 4 different task orders used for our experiments in Table XI.
TABLE IX: The details of 15 datasets in the SuperNI Benchmark [48].
Dataset name Task Metric
1. task639_multi_woz_user_utterance_generation dialogue generation Rouge-L
2. task1590_diplomacy_text_generation dialogue generation Rouge-L
3. task1729_personachat_generate_next dialogue generation Rouge-L
4. task181_outcome_extraction information extraction Rouge-L
5. task748_glucose_reverse_cause_event_detection information extraction Rouge-L
6. task1510_evalution_relation_extraction information extraction Rouge-L
7. task002_quoref_answer_generation question answering Rouge-L
8. task073_commonsenseqa_answer_generation question answering Rouge-L
9. task591_sciq_answer_generation question answering Rouge-L
10. task511_reddit_tifu_long_text_summarization summarization Rouge-L
11. task1290_xsum_summarization summarization Rouge-L
12. task1572_samsum_summary summarization Rouge-L
13. task363_sst2_polarity_classification sentiment analysis accuracy
14. task875_emotion_classification sentiment analysis accuracy
15. task1687_sentiment140_classification sentiment analysis accuracy
TABLE X: The details of 15 classification datasets in the Long Sequence Benchmark [18]. First five tasks correspond to the standard CL benchmark [61].
Dataset name Category Task Domain Metric
1. Yelp CL Benchmark sentiment analysis Yelp reviews accuracy
2. Amazon CL Benchmark sentiment analysis Amazon reviews accuracy
3. DBpedia CL Benchmark topic classification Wikipedia accuracy
4. Yahoo CL Benchmark topic classification Yahoo Q&A accuracy
5. AG News CL Benchmark topic classification news accuracy
6. MNLI GLUE natural language inference various accuracy
7. QQP GLUE paragraph detection Quora accuracy
8. RTE GLUE natural language inference news, Wikipedia accuracy
9. SST-2 GLUE sentiment analysis movie reviews accuracy
10. WiC SuperGLUE word sense disambiguation lexical databases accuracy
11. CB SuperGLUE natural language inference various accuracy
12. COPA SuperGLUE question and answering blogs, encyclopedia accuracy
13. BoolQA SuperGLUE boolean question and answering Wikipedia accuracy
14. MultiRC SuperGLUE question and answering various accuracy
15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
TABLE XI: Four different orders of task sequences used for our experiments. Orders 1-2 correspond to the SuperNI benchmark. Orders 3-4 are long-sequence orders following [18].
Order Model Task Sequence
1 T5, LLaMA-2
task1572 → task363 → task1290 → task181 → task002 →
task1510 → task639 → task1729 → task073 → task1590 →
task748 → task511 → task591 → task1687 → task875
2 T5, LLaMA-2
task748 → task073 → task1590 → task639 → task1572 →
task1687 → task591 → task363 → task1510 → task1729 →
task181 → task511 → task002 → task1290 → task875
3 T5, LLaMA-2
mnli → cb → wic → copa → qqp → boolqa → rte → imdb →
yelp → amazon → sst-2 → dbpedia → ag → multirc → yahoo
4 T5, LLaMA-2
yelp → amazon → mnli → cb → copa → qqp → rte → imdb →
sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic

### v2

@@
 Abstract
-Language model continual learning (CL) has recently garnered significant interest due to its potential to adapt large language models (LLMs) to dynamic real-world environments without re-training.
-A key challenge in this field is catastrophic forgetting, where models lose previously acquired knowledge when learning new tasks.
-Existing methods commonly employ multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge for each task, but these approaches lack efficiency and overlook the potential for knowledge transfer through task interaction.
-In this paper, we present a novel CL framework for language models called Task Skill Localization and Consolidation (TaSL), which enhances knowledge transfer without relying on memory replay.
-TaSL first divides the model into ‘skill units’ based on parameter dependencies, enabling more granular control.
-It then employs a novel group-wise skill localization technique to identify the importance distribution of skill units for a new task.
+Language model continual learning (CL) has recently attracted significant interest for its ability to adapt large language models (LLMs) to dynamic real-world scenarios without retraining.
+A major challenge in this domain is catastrophic forgetting, where models lose previously acquired knowledge upon learning new tasks.
+Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks.
+In this paper, we introduce a novel CL framework for language models, named Task Skill Localization and Consolidation (TaSL), which boosts knowledge transfer without depending on memory replay.
+TaSL initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control.
+Subsequently, it employs a novel group-wise skill localization technique to ascertain the importance distribution of skill units for a new task.
 By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer.
-As a result, TaSL achieves a superior balance between retaining previous knowledge and excelling in new tasks.
-TaSL also shows strong generalizability, suitable for general models and customizable for PEFT methods like LoRA.
-Additionally, it demonstrates notable extensibility, allowing integration with memory replay to further enhance performance.
-Extensive experiments on two CL benchmarks, with varying model sizes (from 220M to 7B), demonstrate the effectiveness of TaSL and its variants across different settings.
+As a result, TaSL achieves an optimal balance between retaining prior knowledge and excelling in new tasks.
+TaSL also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA.
+Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques.
+Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of TaSL and its variants across different settings.
 Index Terms:
-Language model continual learning, catastrophic forgetting, skill localization.
+Language model continual learning, catastrophic forgetting.
 I Introduction
@@
 This capability enables LLMs to dynamically adapt to new tasks and acquire additional knowledge [3, 4].
-An effective CL system must address two critical challenges: (1) Catastrophic Forgetting (CF)  [5], the phenomenon where a model’s proficiency in previous tasks deteriorates as it learns new ones, and (2) Knowledge Transfer (KT)  [6], enhancing task performance through the transfer of knowledge.
+An effective CL system must address two critical challenges: (1) Catastrophic Forgetting (CF)  [5], where a model’s performance on previous tasks deteriorates as it learns new ones, and (2) Knowledge Transfer (KT)  [6], enhancing task performance through the transfer of knowledge.
 KT can be categorized into forward transfer, which improves new task performance using knowledge from previous tasks, and backward transfer, which enhances performance on previous tasks after learning a new relevant task.
@@
 Figure 1: Conceptual illustration of TaSL.
-By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-specific and task-shared parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
-Given the considerable computational demands, recent efforts have explored CL for LLMs using parameter-efficient fine-tuning (PEFT) methods [7, 8], such as Low-Rank Adaptation (LoRA) [7].
-Traditional rehearsal-based approaches, which require storing data examples from past tasks for replay, face significant privacy and memory issues [9, 10, 3].
+By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
+Due to the high computational demands, recent studies have explored CL for LLMs using parameter-efficient fine-tuning (PEFT) methods [7, 8], such as Low-Rank Adaptation (LoRA) [7].
+Traditional rehearsal-based approaches, which require storing data examples from past tasks for replay, face significant privacy and memory issues [9, 3].
 Another prominent line of work in recent studies is the parameter isolation CL methods [10, 11, 12, 13].
-These methods assign a dedicated PEFT block for each new task to acquire task-specific knowledge, preserving the block for selective activation during testing.
+These methods assign a dedicated PEFT block for each new task to capture task-specific knowledge, preserving the block for selective activation during testing.
 However, they exhibit significant limitations in effectively addressing the KT and CF challenges.
-On one hand, the prevalent design of these frameworks, which utilizes multiple PEFT blocks, introduces redundancy and inefficiency [15].
-The number of required PEFT blocks increases with each new task, leading to substantial memory storage costs and making it difficult to handle long task sequences [16, 17].
-Moreover, such approaches essentially create separate expert models for each task, which limits their ability to generalize to unseen tasks [18] (Limitation 1).
-On the other hand, these parameter-isolation CL methods do not consider interactions between tasks, which hinders KT.
-For instance, the method in [12] involves learning each PEFT block separately within individual tasks.
-Similarly, Orthogonal Low-Rank Adaptation (O-LoRA) [19] updates parameters via gradient projection in orthogonal subspaces.
-Although these approaches may mitigate CF to some extent, they cut off the potential transfer of knowledge stored across different PEFT blocks, thus impeding bi-directional KT among various tasks and leading to suboptimal performance (Limitation 2).
+On one hand, the prevalent design of these frameworks, characterized by the use of multiple PEFT blocks, introduces redundancy and operational inefficiency [14].
+As the number of tasks grows, the required PEFT blocks accumulate, leading to increased memory storage demands and making it difficult to handle long task sequences [15, 16].
+Furthermore, such approaches tend to develop separate specialist models for each task, which severely limits their capability to generalize across unseen tasks [17] (Limitation 1).
+On the other hand, these parameter-isolation CL methods overlook the interactions between tasks, which hinders KT.
+For instance, the method in [11] learns each PEFT block separately within individual tasks.
+Similarly, Orthogonal Low-Rank Adaptation (O-LoRA) [18] employs orthogonal subspace gradient projections for parameter updates.
+While these strategies might mitigate CF, they cut off the potential for leveraging knowledge distributed across various PEFT blocks, thus obstructing bi-directional KT among tasks and resulting in less-than-optimal performance (Limitation 2).
 To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay.
 Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20].
-For instance, the authors in [22] uncovered a core region on LLMs crucial for all languages, while distinct monolingual regions exist for different languages.
-Preserving these important regions can help avoid forgetting.
-The research in [15] also found that during fine-tuning with LoRA, many redundant parameter modifications are often retained, suggesting that PEFT blocks contain significant redundancies.
-Based on these insights, TaSL facilitates KT by identifying and consolidating the importance distribution of model parameters across tasks.
-TaSL initially employs a group-wise importance-aware skill localization technique that utilizes gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
+For instance, the authors in [21] uncovered a core region in LLMs that is crucial for all languages, suggesting that preserving these vital regions can mitigate forgetting.
+Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
+Based on these insights, TaSL facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks.
+Initially, TaSL utilizes a group-wise importance-aware skill localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
 By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1.
-Our innovative skill consolidation phase then categorically integrates weights from previous tasks with the current one, enabling effective KT while minimizing forgetting.
+The subsequent skill consolidation phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
 In detail, TaSL first reconstructs the model or PEFT block into fine-grained “skill units”.
@@
 The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task.
-Our approach, focusing on parameter space rather than dataset-driven categorization of task-specific and task-shared knowledge, offers a more effective solution to managing KT in LLMs, overcoming inaccuracies caused by dataset noise.
-Our skill consolidation stage, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge.
-During this phase, we promote forward KT by initializing new tasks with previously fine-tuned weights.
-For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, enhancing their capability.
-To prevent CF, we preserve the integrity of skill units containing previous task-specific knowledge, ensuring they remain unaffected by new task learning (addressing Limitation 2).
+Our skill consolidation phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge.
+This stage facilitates forward KT by using previously fine-tuned weights as the starting point for new tasks.
+For backward KT, we merge knowledge from both current and past tasks into localized task-shared skill units, boosting their effectiveness.
+To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
 Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA.
-While the original TaSL framework has shown promising results, it has only been validated on a specific sub-task, and its effectiveness on general CL benchmarks has yet to be explored.
-Additionally, on the technical side, the skill localization process in TaSL relies on first-order gradients, which may not always provide precise importance localization.
-Furthermore, the skill consolidation phase in TaSL involves many hyperparameters, potentially complicating the averaging process.
-To address these issues, we propose TasLoRA, a LoRA-tailored version of TaSL.
-TasLoRA reconstructs the LoRA adapter into new skill units based on parameter dependencies, ensuring more efficient management of knowledge through sequential task learning.
-During the fine-tuning process, an additional orthogonal loss is added to ensure that skill units within the same LoRA adapter capture distinct latent semantic features.
-For skill localization, TasLoRA employs a novel second-order gradient approximate group-wise metric to compute importance scores more precisely.
-For skill consolidation, we transition from the previous hard-mask model averaging approach to a soft-masking strategy, using an adaptive averaging technique.
-This allows for a more flexible integration of task-specific and shared parameters, dynamically adjusting the contribution of each skill unit based on its importance.
-Furthermore, to adapt to a broader range of application scenarios, we combine TaSL with memory replay, introducing the TaSL-M model.
-Through extensive experiments on two CL benchmarks with different parameter-level backbones (from 220M to 7B), our TaSL framework, along with its variants, excels in mitigating CF and showcases remarkable capabilities for KT, outperforming state-of-the-art (SOTA) methods.
+While the initial TaSL framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored.
+Additionally, the reliance on first-order gradients for skill localization in TaSL may not yield precise importance localization.
+The skill consolidation phase also involves many hyperparameters, which could complicate the averaging process.
+To address these issues, we propose TasLoRA, a variant of TaSL optimized for LoRA.
+TasLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning.
+It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter.
+For skill localization, TasLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores.
+In skill consolidation, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit.
+Additionally, we combine TaSL with memory replay to broaden its application scope, introducing the TaSL-M model.
+Through extensive experiments on two CL benchmarks with four parameter-level backbones, our TaSL framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
 The main contributions are summarized as follows:
-  • We propose a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning.
-By identifying and consolidating task-specific and task-shared knowledge at a granular skill unit level, TaSL achieves effective knowledge transfer and mitigates catastrophic forgetting, overcoming the limitations of previous approaches.
-  • We develop various group-wise skill localization and fine-grained skill consolidation techniques.
-For instance, the parameter importance metric in skill localization can be based on first-order gradients or a new second-order gradient approximation method.
-In skill consolidation, our model averaging strategies include a categorical hard-mask approach and an adaptive soft-mask method.
-  • The TaSL framework demonstrates strong generalizability and extensibility.
-The flexible design of skill units enables TaSL to be easily tailored to PEFT methods, such as LoRA, optimizing performance for specific model architectures.
-Additionally, TaSL can be integrated with memory replay technique to further enhance performance and adapt to a broader range of application scenarios.
-  • Extensive evaluation on two CL benchmarks underscores the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
-Furthermore, TaSL consistently excels across diverse model sizes (ranging from 220M to 7B), different model architectures (T5 and LLaMA-2), and unseen tasks.
+  • We introduce the Task Skill Localization and Consolidation (TaSL) framework for language model continual learning.
+By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, TaSL ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
+  • We devise various group-wise skill localization and fine-grained skill consolidation techniques.
+Notably, our skill localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance.
+For skill consolidation, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
+  • The TaSL framework is distinguished by its robust generalizability and extensibility.
+Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures.
+Moreover, integrating TaSL with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
+  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
+Furthermore, TaSL consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
 II Related Work
-II-A Continual Learning
-Continual learning [3] aims to develop learning algorithms capable of accumulating knowledge from non-stationary data.
-Conventional Continual Learning methods can be divided into three categories: (i) Regularization-based methods add explicit regularization terms to preserve the knowledge of previous tasks [13, 23].
-(ii) Rehearsal-based methods address catastrophic forgetting by storing old training samples [24, 25] or by training generative models to provide pseudo samples of previous tasks [26, 27].
-Due to the memory and data privacy issues associated with these methods, we focus on rehearsal-free CL methods.
-(iii) Parameter isolation-based methods dynamically expand model capacity or assign isolated parameters dedicated to each task throughout the CL process to prevent interference between tasks [28, 29, 30].
-Language Model Continual Learning with PEFT.
-Based on PEFT methods, current approaches for the continual learning of LLMs adopt the concept of parameter isolation, using a pipeline fashion to learn and select PEFT blocks for each task [12, 31, 32].
-However, these approaches are limited in effectively addressing the challenges of catastrophic forgetting (CF) and knowledge transfer (KT) [33].
-For instance, the authors in [19] and [34] constrain the learning of PEFT blocks to maintain orthogonality, restricting KT among different tasks.
-Although the recent SAPT method [16] and DAP method [35] achieves KT, it relies on pseudo sample replay or unlabeled domain corpora.
-Our method achieves KT without memory replay, offering unique advantages in data privacy and parameter efficiency.
-By using layering low-rank adapters on the key and value projection matrices of transformer blocks, we balance retaining previous knowledge and excelling in new tasks.
+II-A Language Model Continual Learning
+Continual learning [3] seeks to develop algorithms that accumulate knowledge from non-stationary data sources.
+Conventional continual learning methods can be categorized into three main types: (i) Regularization-based methods impose explicit constraints to preserve the knowledge of previous tasks by penalizing the variation of each parameter based on its contribution to past tasks [12, 22], such as in the EWC [23] method.
+However, these methods typically restrict the learning of task-shared knowledge by only considering parameter importance from the perspective of past tasks.
+This limitation often leads to a suboptimal balance between retaining previous knowledge and excelling at new tasks, as it does not account for the relevance of parameters to both past and current tasks.
+(ii) Rehearsal-based methods mitigate catastrophic forgetting by either storing old training samples [24, 25] or training generative models to provide pseudo-samples of previous tasks [26, 27], enabling data replay during the learning of new tasks.
+(iii) Architecture-based methods aim to prevent task interference by dynamically expanding model capacity or assigning dedicated parameter subspaces to each task [28, 29].
+While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks.
+Our TaSL framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting.
+To further enhance performance, we extend TaSL with memory replay techniques, resulting in TaSL-M, improving its flexibility and applicability across various scenarios.
+Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy.
+This approach involves using a pipeline to learn and select PEFT blocks for each task [11, 30].
+However, as the number of tasks increases, the accumulation of PEFT blocks can become inefficient, and achieving KT between blocks remains challenging.
+More recently, authors in [31] have explored using the Mixture of Experts (MoE) approach to connect multiple PEFT blocks, but this essentially relies on coarse-grained model averaging.
+This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting.
+To address these challenges in the context of PEFT-based CL, we extend the TaSL framework and introduce TasLoRA.
+By dividing fine-grained skill units based on parameter dependencies within LoRA, TasLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
 II-B Skill Localization
-Research shows that model parameters do not all contribute equally to performance [36].
-The authors in [21] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, suggesting that fine-tuning only these critical parameters can achieve results comparable to full model fine-tuning.
-However, their method involves additional steps to identify and retrain these key parameters during the post-fine-tuning phase, which reduces efficiency.
-Inspired by techniques from the pruning community, previous studies have utilized gradient-based metrics to pinpoint important parameters during fine-tuning.
-Sensitivity-based scoring [37, 38] evaluates the impact on training loss, and sensitivity smoothing, as employed by [39], helps remove unnecessary parameters for more efficient fine-tuning.
-Nonetheless, these methods often result in element-wise pruning, where individual parameters are removed based on their importance scores.
-This approach can be computationally and storage-intensive, as it requires detailed management of numerous parameters throughout the model.
-Building on these advancements, we propose a novel importance-aware skill localization method.
-This approach uniquely differentiates between task-specific and shared parameters to mitigate forgetting and achieve KT in CL.
+Research has shown that model parameters contribute unevenly to performance [32].
+For example, the authors in [21] identified a core region in LLMs linked to linguistic competence; freezing this region during further pre-training has been shown to mitigate catastrophic forgetting.
+Additionally, the authors in [33] found neurons within LLMs that store personally identifiable information, highlighting the potential for improved privacy protections.
+However, these studies primarily focus on describing such findings rather than applying them to solve practical problems.
+Motivated by these findings, we address the challenge of catastrophic forgetting in CL.
+While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency.
+In contrast, our importance-aware skill localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time.
+Combined with our skill consolidation strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
 III Preliminaries
-III-A Continual Learning
-Continual learning [40] aims to create learning algorithms that can continuously accumulate knowledge from ongoing sequences.
-In supervised continual learning, a sequence of tasks $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{K}\}$ is presented in a streaming manner.
+III-A Continual Learning Setup
+Continual learning [2] aims to develop algorithms that can progressively accumulate knowledge from ongoing sequences.
+In supervised continual learning, a series of tasks $\{\mathcal{T}_{1},\ldots,\mathcal{T}_{K}\}$ is presented sequentially.
 Each task $\mathcal{T}_{k}$ includes a distinct target dataset $\mathcal{D}_{k}=\left\{\left(x_{i}^{k},y_{i}^{k}\right)\right\}_{i=1}^{N_{k}}$ of size $N_{k}$, where $x_{i}^{k}\in\mathcal{X}_{k}$ and $y_{i}^{k}\in\mathcal{Y}_{k}$.
-The model must adapt to these tasks sequentially, having access only to $\mathcal{D}_{k}$ during the $k$-th task.
-Generally, with a prediction model $f$ parameterized by $\Theta$, continual learning aims to optimize the following objective across all tasks:
+The model, parameterized by $\Theta$, is trained to adapt to these tasks one at a time, accessing only $\mathcal{D}_{k}$ during the $k$-th task.
+The overarching goal of continual learning is to optimize the following objective:
 $\max_{\Theta}\sum_{k=1}^{K}\sum_{x,y\in\mathcal{D}_{k}}\log p_{\Theta}(y\mid x)$ (1)
-The notation $f_{k}$ refers to the model after training on task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ denotes the model after averaging for $\hat{f}_{k-1}$ and $f_{k}$.
-Our TaSL framework aims to address a more challenging scenario where the model cannot access any historical data during training [19].
-Nevertheless, to adapt to a wider range of application scenarios, we have also combined TaSL with memory replay.
-In this setting, we randomly save $\left|\mathcal{M}\right|$ samples from the training set of each previous task $\mathcal{T}_{i}$ in memory $\mathcal{M}_{i}$ and jointly train the model on new task data $\mathcal{D}_{k}$ and memory $\mathcal{M}_{<k}$.
-This extension, called TaSL-M, leverages memory replay to enhance performance.
+And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging.
+Our TaSL framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18].
+To broaden its applicability, we have augmented TaSL with memory replay.
+In this enhanced setup, known as TaSL-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$.
+The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.
 III-B Low-Rank Adaption
 Low-Rank Adaption (LoRA) [7] is a parameter-efficient fine-tuning method to adapt LLMs to novel tasks.
-It assumes that the changes of parameters lie in a low-rank space when the model is fully fine-tuned on a downstream task.
-This approach eliminates the necessity of fine-tuning the entire model, thereby improving computational efficiency and resource utilization.
+It operates under the assumption that parameter changes during full fine-tuning on a downstream task reside within a low-rank space.
+This approach allows for not having to fine-tune the entire model, thus significantly improving computational efficiency and resource utilization.
 Inspired by the low-rank internal dimensionality [34], LoRA hypothesizes the updates to the weights also has a low “intrinsic rank” during adaptation.
-For a pre-trained weight matrix $W^{(0)}\in\mathbb{R}^{out\times in}$ that takes $x$ as input, LoRA modifies the output $h$ of $W^{(0)}$ with a low-rank decomposition:
+For instance, consider a pre-trained weight matrix $W^{(0)}\in\mathbb{R}^{out\times in}$ that takes $x$ as input, LoRA modifies the output $h$ of $W^{(0)}$ with a low-rank decomposition:
 $h=W^{(0)}x+\Delta Wx=W^{(0)}x+BAx,$ (2)
@@
 $A$ is initialized from a random Gaussian distribution and $B$ is initialized with all zeros.
-During the training stage, $W^{(0)}$ is fixed, and only $BA$ is updated.
-Due to its superior performance, LoRA has become one of the most popular PEFT methods in NLP community.
+During training, $W^{(0)}$ remains fixed, and only $BA$ is updated.
+Due to its effectiveness, LoRA has emerged as one of the most favored PEFT methods within the NLP community [35].
 IV Proposed Method: TaSL
 IV-A Overview
-TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to accurately identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs a novel fine-grained model averaging strategy to integrate model weights from both current and past tasks for effective knowledge transfer.
-Figure 2 provides a comprehensive overview of our proposed TaSL framework, with the following subsections detailing each component and the corresponding extensions.
-Figure 2: Left: Illustration of reconstructing LoRA as fine-grained skill units.
+TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
+Figure 2 presents a detailed overview of our proposed TaSL framework, with the following subsections elaborating on each component and the corresponding extensions.
+Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units.
 Right: Overview of TaSL.
@@
 Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$.
-The integration is guided by the importance distributions of skill units across various tasks.
-This process can be repeated with the introduction of each subsequent task.
+This integration is strategically guided by the importance distributions of skill units across different tasks.
+This iterative process continues with the addition of each new task.
 IV-B Fine-Grained Skill Unit
-Before we localize the importance of parameters, we need to define a structure that helps us better organize the knowledge stored in the model.
-This structure will provide a clearer understanding of how to mitigate CF and achieve KT.
-Therefore, we refer to the basic unit of stored skills or knowledge in the model as a “skill unit.” Based on different usage scenarios, we offer the following two structures for skill units:
+Before localizing the importance of parameters, it is essential to define a structure that facilitates better organization and understanding of the knowledge stored within the model.
+This structure aids in addressing CF and enhancing KT.
+We define this basic structural element of stored skills or knowledge in the model as a “skill unit.” Depending on different usage scenarios, we propose two structures for skill units:
 IV-B1 Matrix-Level Skill Unit
 In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA.
-This approach aligns with the original TaSL framework.
-The advantage of this division is that it is model-agnostic, making it applicable to both traditional full-parameter fine-tuning and the latest parameter-efficient fine-tuning methods, such as LoRA.
-This provides a higher degree of generalizability.
+This approach aligns with the original TaSL framework and offers the advantage of being model-agnostic.
+It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
 IV-B2 LoRA-Tailored Skill Unit
-While the matrix-level division is simple and intuitive, it did not sufficiently consider the intra-matrix redundancy and inter-matrix dependencies.
-To address these limitations and the inefficiencies noted in recent work [19], which treats each LoRA adapter as a container for task-specific knowledge, we decompose the model matrices into new, finer-grained “skill units” based on parameter dependencies tailored for LoRA.
-The construction process for these LoRA-tailored skill units is as follows.
+While the matrix-level division is straightforward and intuitive, it does not adequately address intra-matrix redundancy or inter-matrix dependencies.
+To address these limitations and the inefficiencies noted in recent work [18], which treats each LoRA adapter as a container for task-specific knowledge, we propose a more refined decomposition of model matrices into new, finer-grained “skill units” based on parameter dependencies specifically tailored for LoRA.
+The construction process for these LoRA-tailored skill units unfolds as follows.
 As shown in Eq (2), $A$ and $B$ can be viewed as a combination of a set of vectors: $A=[a_{1},a_{2},\cdots,a_{r}]$, $B=[b_{1},b_{2},\cdots,b_{r}]$, where $a_{i}\in\mathbb{R}^{in}$, $b_{i}\in\mathbb{R}^{out}$.
@@
 $=W^{(0)}+u_{1}+u_{2}+\cdots+u_{r},$
-where skill unit $u_{i}$ is a matrix obtained by the product of the vectors $b_{i},a_{i}$. Therefore, LoRA can be viewed as a fusion of knowledge learned from multiple skill units:
+where each skill unit $u_{i}$ is a matrix formed by the product of the vectors $b_{i},a_{i}$. Thus, LoRA can be conceptualized as an aggregation of knowledge across multiple skill units:
 $W$ $=W^{(0)}+\sum_{i=1}^{r}b_{i}a_{i}=W^{(0)}+\sum_{i=1}^{r}u_{i}$ (4)
-This division strategy effectively decomposes the parameters of a single LoRA adapter, reducing the redundancy in the traditional approach of employing separate LoRA adapters for each task [19].
-To ensure that different skill units within the same layer learn distinct latent semantic features, we apply a regularization term as introduced by [39]:
+This division strategy significantly reduces the redundancy inherent in the traditional approach of utilizing separate LoRA adapters for each task [18].
+To ensure that different skill units within the same layer learn distinct latent semantic features, we incorporate a regularization term as introduced in [36]:
 $R(A,B)=\|A^{T}A-I\|_{F}^{2}+\|B^{T}B-I\|_{F}^{2},$ (5)
-where $I$ is an identity matrix, this forces $A$ and $B$ to be orthogonal after training.
+where $I$ is the identity matrix, this forces $A$ and $B$ to be orthogonal after training.
 IV-C Importance-aware Skill Localization
-To calculate the importance of each skill unit $u$, we introduce a group-wise metric that also addresses the significant computational and storage burdens associated with previous parameter-level importance calculation methods [42]:
+To calculate the importance of each skill unit $u$, we introduce a group-wise metric that mitigates the significant computational and storage burdens associated with previous parameter-level importance calculation methods [37]:
 $\mathcal{I}(u)=\frac{1}{in\times out}\sum\limits_{i=1}^{in}\sum\limits_{j=1}^{out}s(w_{ij})$ (6)
 where $w_{ij}$ denotes the trainable parameters, and $in\times out$ represents the total parameter count in a skill unit $u$.
-$\mathcal{I}(u)$ measures the collective importance of all parameters within each skill unit, where higher values signify increased importance.
+The function $\mathcal{I}(u)$ measures the collective importance of all parameters within each skill unit, with higher values indicating greater importance.
 The importance function $s(\cdot)$ for individual parameters, inspired by the pruning community [38], is quantified by measuring the impact of its removal on the loss.
-Specifically, to estimate the importance of $W_{i}$, we calculate the change in loss when the parameter is zeroed out using:
+Specifically, to estimate the importance of $W_{i}$, we measure the change in loss when the parameter is zeroed out using:
 $I_{W_{i}}=|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}_{W_{i}}(\mathcal{D})-\mathcal{L}_{W_{i}=0}(\mathcal{D})|$ (7)
 $=\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W_{i}}W_{i}-\frac{1}{2}W_{i}H_{ii}W_{i}+\mathcal{O}(\|W_{i}\|^{3})\right|$
-where $H$ is the hessian matrix, $W_{i}$ represents the $i$-th parameter in $W$, and $\mathcal{L}$ represents the next-token prediction loss.
-If removing a parameter has a significant influence, then the model is sensitive to it, and we should retain it [44].
+where $H$ is the Hessian matrix, $W_{i}$ is the $i$-th parameter in $W$, and $\mathcal{L}$ is the next-token prediction loss.
+If removing a parameter has a significant influence, then the model is sensitive to it, and we should retain it.
 Based on Eq.
-(7), we can derive the following two metrics for calculating parameter importance:
+(7), we can derive the following two metrics for calculating importance:
 IV-C1 First-Order Gradient-Based Metric
-In original TaSL framework, we followed the standard approach, defining $s(\cdot)$ as the magnitude of the gradient-weight product:
+In the original TaSL framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
 $I_{W_{i}}=\left|W_{i}\nabla_{W_{i}}\mathcal{L}\right|$ (8)
-This metric only considers the first-order gradient term in the importance calculation.
-However, the importance of a parameter is also reflected in the second-order gradients (Hessian matrix).
-Ignoring second-order gradient information may lead to biases in localization, thus affecting model performance.
+This metric primarily considers the first-order gradient term in determining parameter importance.
+However, the significance of a parameter is also influenced by second-order gradients, represented by the Hessian matrix.
+Overlooking this second-order gradient information can introduce biases in the localization process, potentially degrading model performance.
 Therefore, we propose the following new metric.
 IV-C2 Second-Order Gradient Approximation Metric
-Direct computation of the Hessian matrix for LLMs is impractical due to its $\mathcal{O}(N^{2})$ complexity.
-To address this, we propose a second-order gradient approximation for the importance metric, which balances the efficiency of first-order methods with the precision of second-order gradients.
-To reduce the computational complexity, the diagonal of the Hessian $H_{ii}$ can be approximated by the Fisher information matrix [45], and the importance can be defined as:
+The direct computation of the Hessian matrix for LLMs is impractical due to its $\mathcal{O}(N^{2})$ complexity.
+To address this, we propose a second-order gradient approximation metric that balances the efficiency of first-order methods with the precision of second-order gradients.
+To reduce computational demands, we approximate the diagonal of the Hessian, $H_{ii}$, using the Fisher information matrix [39].
+The importance is then defined as:
 $I_{W_{i}}\approx\left|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial W_{i}}W_{i}-\frac{1}{2}\sum_{j=1}^{N}\left(\frac{\partial\mathcal{L}(\mathcal{D}_{j})}{\partial W_{i}}W_{i}\right)^{2}\right|$ (9)
 However, calculating the importance as specified in Eq.
-(9) across the entire training set introduces significant challenges, as model training typically only has access to mini-batch data.
-This means the metric is subject to variability due to stochastic sampling and training dynamics, introducing large uncertainty in estimating sensitivity.
-To mitigate this, we propose a more reliably importance metric based on sensitivity smoothing and uncertainty quantification [39]:
+(9) across the entire training set introduces significant challenges, as model training typically accesses only mini-batch data.
+This limitation subjects the metric to variability due to stochastic sampling and training dynamics, introducing substantial uncertainty in estimating parameter sensitivity.
+To mitigate this, we propose a more reliable importance metric based on sensitivity smoothing and uncertainty quantification [36]:
 $\begin{split}\bar{I}^{(t)}\left(w_{ij}\right)=\alpha_{1}\bar{I}^{(t-1)}\left(w_{ij}\right)+\left(1-\alpha_{1}\right)I^{(t)}\left(w_{ij}\right)\end{split}$ (10)
@@
 $\bar{I}^{(t)}$ represents smoothed sensitivity and $\bar{U}^{(t)}$ is the uncertainty term quantified by the local variation between $I^{(t)}$ and $\bar{I}^{(t)}$.
-Using the exponential moving average of the importance metric, we can retain and explore the trajectory gradient for a longer time, providing a more robust and precise importance assessment.
-Importance is then determined by:
+Using the exponential moving average of the importance metric, we can retain and examine the trajectory gradient for a longer time, yielding a more precise importance assessment.
+Importance is ultimately determined by:
 $s^{(t)}\left(w_{ij}\right)=\bar{I}^{(t)}\left(w_{ij}\right)\cdot\bar{U}^{(t)}\left(w_{ij}\right)$ (12)
-To compute the importance score of each skill unit for current task $\mathcal{T}_{k}$, we employ Eq.
+To calculate the importance score of each skill unit for the current task $\mathcal{T}_{k}$, we use Eq.
 (6) during fine-tuning.
@@
   $f_{k}$ and importance scores $\mathcal{I}(\mathcal{U}_{k})$ for $\mathcal{U}_{k}$.
-After computing importance scores for each skill unit at the current task $\mathcal{T}_{k}$, it is essential to compare these with scores from all previously learned tasks to distinguish between task-specific and task-shared skill units.
-To avoid the inefficiency of storing scores for each past task, we aggregate importance scores from all prior tasks into a cumulative score for tasks up to $\mathcal{T}_{k-1}$.
-This method allows for the iterative refinement of accumulated scores without separately saving past task scores.
+After computing the importance scores for each skill unit at the current task $\mathcal{T}_{k}$, it is crucial to compare these with scores from all previously learned tasks to distinguish between task-specific and task-shared skill units.
+To streamline this process and avoid the inefficiencies associated with storing scores for each past task, we aggregate importance scores from all prior tasks into a cumulative score for tasks up to $\mathcal{T}_{k-1}$.
+This approach allows for iterative refinement of accumulated scores without the need to separately store past task scores.
 The skill units with these cumulative scores up to $\mathcal{T}_{k-1}$ are denoted as $\hat{\mathcal{U}}_{k-1}$, calculated using:
 $\begin{split}\mathcal{I}(\hat{\mathcal{U}}_{k-1})=\beta\text{Norm}(\mathcal{I}(\hat{\mathcal{U}}_{k-2}))+(1-\beta)\text{Norm}(\mathcal{I}(\mathcal{U}_{k-1}))\end{split}$ (13)
-where $\beta\in[0,1]$, and Norm($\cdot$) normalizes importance scores to the [0, 1] range, thus resolving discrepancies across models.
-The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set to be equal to $\mathcal{I}(\mathcal{U}_{1})$.
+where $\beta\in[0,1]$ and Norm($\cdot$) normalizes the importance scores to the [0, 1] range, ensuring consistency across models.
+The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set equal to $\mathcal{I}(\mathcal{U}_{1})$.
 Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
 IV-D Skill Consolidation
-After skill localization, the subsequent vital phase involves consolidating the knowledge in the skill unit into a unified framework.
-This process demands a sophisticated model averaging approach considering various factors to optimize task performance.
-Traditional coarse-grained model averaging assumes that all model weights are equally important for the training task [46, 47], which can be written as the following iterative computation format:
+Following skill localization, the next critical phase is skill consolidation, where knowledge within each skill unit is integrated into a cohesive framework.
+This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
+Traditional coarse-grained model averaging operates under the assumption that all model weights hold equal importance for the training task [23, 40], typically implemented through the following iterative computation:
 $\hat{f}_{k}=\lambda\hat{f}_{k-1}+\left(1-\lambda\right)f_{k}$ (14)
 However, coarse-grained methods may overemphasize weights that are irrelevant to the current task, contaminating previously acquired task-specific knowledge and leading to forgetting.
-To address this issue, we introduce two fine-grained averaging strategies that focus on skill units rather than the entire model.
-Our approach distinguishes between task-shared and task-specific skill units, applying weighted averaging to parameters within each unit.
-This tailored method provides a customized solution to meet the distinct needs of each task.
+To overcome these limitations, we introduce two fine-grained averaging strategies that focus on skill units instead of the entire model.
+Our method distinguishes between task-shared and task-specific skill units, applying a weighted averaging technique to the parameters within each unit.
+This refined approach offers a tailored solution, addressing the unique requirements of each task more effectively.
 IV-D1 Static Weighted Consolidation
-This strategy, used in the original TaSL framework, begins by setting importance thresholds $\delta$ using quantiles to select the top 20% of skill units based on importance scores.
-A skill unit $u_{i}^{k}$ is deemed important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ is above $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
-Our static weighted consolidation strategy customizes parameter combination for each skill unit, based on its importance under different tasks, as follows:
+In the original TaSL framework, this consolidation strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores.
+A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
+The static weighted consolidation strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows:
 $\hat{u}_{i}^{k}=\begin{cases}\gamma\hat{u}_{i}^{k-1}+(1-\gamma)u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{+}$}\\ \hat{u}_{i}^{k-1},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{-}$}\\ u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{+}$}\\ \frac{1}{2}(\hat{u}_{i}^{k-1}+u_{i}^{k}),&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{-}$}\\ \end{cases}$ (15)
-This approach performs the element-wise adjustment of parameters within each skill unit based on its relevance to previous and current tasks, using hyperparameter $\gamma$ to control their influences.
+This approach performs the element-wise adjustment of parameters within each skill unit based on its relevance to previous and current tasks.
+The hyperparameter $\gamma$ is employed to control their influences.
 IV-D2 Adaptive Weighted Consolidation
-Building upon the original TaSL framework, we propose an Adaptive Weighted Consolidation strategy as our latest extension.
-In contrast to the manual setting of multiple hyperparameters required in the original method, this adaptive strategy simplifies the process by automatically adjusting to various scenarios, making the averaging process more efficient and flexible.
-In detail, for a specific skill unit $u_{i}$, the weighting coefficients are:
+As a sophisticated extension to the static method, which necessitated the manual setting of multiple hyperparameters, this adaptive strategy simplifies the process by automatically adjusting to various scenarios.
+This enhancement increases the efficiency and flexibility of the averaging process.
+Specifically, for a specific skill unit $u_{i}$, the weighting coefficients are determined by:
 $\begin{split}\hat{\lambda}^{k-1}_{i}=\exp\left(\mathcal{I}(\hat{u}^{k-1}_{i})/\tau\right),\lambda^{k}_{i}=\exp\left(\mathcal{I}(u^{k}_{i})/\tau\right)\end{split}$ (16)
@@
 $\begin{split}\hat{u}_{i}^{k}=\left(\frac{\hat{\lambda}^{k-1}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot\hat{u}_{i}^{k-1}+\left(\frac{\lambda^{k}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot u_{i}^{k}\end{split}$ (17)
-By using Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
-  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high while $\mathcal{I}(u^{k}_{i})$ is low, it indicates that the skill unit $u_{i}$ is more important for the history task.
+By applying Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
+  • Mitigating CF: When $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high and $\mathcal{I}(u^{k}_{i})$ is low, this indicates that the skill unit $u_{i}$ is more critical for previous tasks.
 According to Eq.
-(15) or (17), we maintain the previous task-specific knowledge stored in $\hat{u}^{k-1}_{i}$ untouched to prevent the contamination of historical knowledge with task-irrelevant information in $u^{k}_{i}$, thereby mitigating CF.
-  • If $\mathcal{I}(\hat{u}^{k-1}_{i})$ is low and $\mathcal{I}(u^{k}_{i})$ is high, it suggests that the skill unit $u_{i}$ is more important for the current task.
-Since the model’s training initializes from the endpoint of the previous task, the historically learned knowledge is utilized to enhance performance on the current task.
-So we ensure the integrity of the parameters in $u^{k}_{i}$, which stores current task-specific knowledge, thereby facilitating forward KT.
-  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are high, it indicates that the skill unit $u_{i}$ is important for both the previous and current tasks.
-Consequently, we integrate newly acquired knowledge into this task-shared skill unit to enable backward KT.
-  • If both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are low, it indicates the skill unit is not significantly relevant to any task.
-In this case, simply averaging the parameters within this unit is sufficient.
-Skill consolidation is performed before starting a new task in CL, utilizing the averaged model for subsequent task initialization.
-Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are retained for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$.
+(15) or (17), we maintain the previous task-specific knowledge in $\hat{u}^{k-1}_{i}$ untouched, preventing contamination from the less relevant $u^{k}_{i}$.
+  • Facilitating Forward KT: Conversely, if $\mathcal{I}(\hat{u}^{k-1}_{i})$ is low and $\mathcal{I}(u^{k}_{i})$ is high, it suggests that $u_{i}$ holds greater importance for the current task.
+As training initiates from the endpoint of the previous task, maintaining the integrity of parameters in $u^{k}_{i}$ leverages historically learned knowledge to enhance performance on the current task.
+  • Enabling Backward KT: When both $\mathcal{I}(\hat{u}^{k-1}_{i})$ and $\mathcal{I}(u^{k}_{i})$ are high, $u_{i}$ is vital for both previous and current tasks.
+Here, we integrate newly acquired knowledge into this task-shared skill unit to support backward KT.
+  • Handling Low Relevance: If both $\mathcal{I}(\hat{u}^{k-1}{i})$ and $\mathcal{I}(u^{k}{i})$ are low, indicating minimal relevance to any task, a simple averaging of parameters within this unit is adequate.
+Skill consolidation is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization.
+Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$.
 Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
@@
 V TaSL with Memory Replay: TaSL-M
-To adapt TaSL to scenarios where historical data can be used, we propose a memory replay-enhanced version, TaSL-M, to further improve model performance.
+To adapt TaSL for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, TaSL-M, designed to further improve model performance.
 The implementation process is shown in Figure 3.
-Figure 3: Overview of TaSL-M.
-After training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$.
-Before the next task arrives, we introduce a replay stage, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$.
+Figure 3: Overview of TaSL-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of TaSL.
+Following the training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$.
+Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$.
 This process results in a new model $\hat{f}^{m}_{k}$, which helps recover forgotten knowledge.
-Specifically, during the replay stage, we do not fine-tune the parameters of all skill units using historical data, as this could negatively impact the performance on the current task.
-Instead, we fix the current task-specific skill units and fine-tune only the remaining parameters.
+Specifically, during the replay stage, we avoid fine-tuning the parameters of all skill units with historical data, as this could potentially degrade the model’s performance on the current task.
+Instead, we maintain the integrity of the current task-specific skill units and selectively fine-tune the remaining parameters.
 This approach achieves a better balance between retaining previous knowledge and excelling in new tasks.
@@
 For each task, 1,000 instances from the dataset are randomly selected for training, and 100 instances are used for validation and testing.
-Additionally, we employ the Long Sequence Benchmark [18], a continual learning benchmark consisting of 15 classification datasets.
-In line with [19], we randomly select 1,000 samples for training each task and reserve 500 samples per class for validation and testing.
-We explore two different task orders for each benchmark.
-Please refer to the Appendix for more details about the tasks and orders.
-TABLE I: The overall results on two continual learning benchmarks with T5-large model.
-All results are averaged over two different orders of each benchmark.
-Our rehearsal-free TaSL outperforms the previous best method, O-LoRA, by achieving an average 2.3% absolute improvement on AP, an average 2.8% absolute improvement on FWT, and an 4.8% absolute increase in BWT.
+Additionally, we utilize the Long Sequence Benchmark [17], which consists of 15 classification datasets tailored for CL.
+Consistent with previous work [18], we randomly select 1,000 samples per task for training and reserve 500 samples per class for validation and testing.
+The experiments include two different task orders for each benchmark.
+Further details about the tasks and their order are provided in the Appendix.
+TABLE I: The overall results on two continual learning benchmarks with T5-large model. All results are averaged over two different orders of each benchmark.
 Method Memory-Replay SuperNI Benchmark Long Sequence Benchmark
@@
 VI-A3 Baselines
-We evaluate TaSL against the following PEFT-based continual learning baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
+We evaluate TaSL against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
 IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks.
 Replay: replays real samples from old tasks when learning new tasks to avoid forgetting.
-EWC [46]: finetune the model with a regularization loss that prevents updating parameters that could interfere with previously learned tasks.
+EWC [23]: finetune the model with a regularization loss designed to minimize updates to parameters critical to previously learned tasks.
 L2P [42]: uses the input to dynamically select and update prompts from a fixed prompt pool.
-LFPT5 [50]: continuously trains a soft prompt for each task with generative replay and an auxiliary loss.
+LFPT5 [43]: continuously trains a soft prompt for each task with generative replay.
 Prog-Prompt [17]: sequentially concatenates previous learned prompts to the current one during the training and testing time.
-O-LoRA [19]: learns tasks in different LoRA subspaces that are kept orthogonal to each other and sums all LoRA weights up at testing time.
+O-LoRA [18]: learns tasks in distinct LoRA subspaces, ensuring orthogonality, and aggregates all LoRA weights for testing.
 SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
-Table II compares our TaSL with these baselines, highlighting three distinct advantages: data privacy-friendliness, model parameter-friendliness, and generalization-friendliness.
-To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our base framework utilizes original components, including matrix-level skill units, first-order gradient-based metric for skill localization (Eq.
-8), and static weighted consolidation (Eq.
+Table II compares our TaSL with these baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities.
+To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for skill localization (Eq.
+8), and employs static weighted consolidation (Eq.
 15).
-TasLoRA: as the latest extension of TaSL, it employs LoRA-tailored skill units, the second-order gradient approximation metric for skill localization (Eq.
-9), and adaptive weighted consolidation (Eq.
+TasLoRA: as the most recent extension of TaSL, this variant introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric for skill localization (Eq.
+9), and features adaptive weighted consolidation (Eq.
 17).
-TasL-M and TasLoRA-M: these variants combine TaSL or TasLoRA with memory replay techniques to further enhance performance.
+TasL-M and TasLoRA-M: these enhancements of TaSL and TasLoRA integrate memory replay to further boost performance.
 TABLE II: The comparison between TaSL and other CL methods.
@@
 VI-A4 Implementation Details
-We utilize two language models adopted by the previous lines of works in CL for NLP: encoder-decoder T5 model [57] and decoder-only LLaMA model [58].
+We utilize two distinct language model architectures previously adopted in the field of CL for NLP: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50].
 In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq.
@@
 (16) to 0.15.
-Following [59], the volume of replay samples is set to 2% of the original training set for all replay-based baseline methods.
+Following [51], we allocate 2% of the original training set volume for replay samples in all replay-based baseline methods.
 For different backbones, we utilized the following hyperparameters:
-  • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, rank = 8, targeting modules [q, v] and 10 epochs.
-  • LLaMA (7B): With a learning rate of 3e-4, batch size of 128, a cutoff length of 512, and 10 epochs.
+  • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, a rank of 8, targeting modules [q, v] and 10 epochs.
+  • LLaMA (7B): With a learning rate of 3e-4, a batch size of 128, a cutoff length of 512, and 10 epochs.
 LoRA settings were rank = 8, alpha = 16, dropout = 0.05, targeting modules [q_proj, v_proj].
 For testing, settings included temperature = 0.02, top_p = 0, top_k = 1, max new tokens = 128.
-Experiments are carried out using 4 NVIDIA A100.
+Experiments are conducted using 4 NVIDIA A100 GPUs.
 VI-B Main Results
-Overall CL results of different methods at the same T5-large backbone are summarized in Table I.
-Firstly, TaSL and its variants demonstrate superior CL performance through effective knowledge transfer.
-Compared to previous rehearsal-free methods such as IncLoRA and O-LoRA, TaSL excels in addressing the two critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
-Moreover, our enhanced version, TasLoRA, shows significant improvements over TaSL.
-Specifically, TasLoRA achieves average gains of 1.3% on AP and 1.4% on BWT, indicating that our extensions to each component are effective.
-Notably, even without relying on memory replay, our method nearly matches the performance of SAPT with memory replay on Long Sequence Benchmark, particularly in BWT, with only a slight difference of 0.8% (-2.04% vs.
+The overall CL results of various methods using the same T5-large backbone are summarized in Table I.
+Firstly, TaSL and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer.
+Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
+Furthermore, our enhanced version, TasLoRA, significantly outperforms TaSL.
+Specifically, TasLoRA achieves average gains of 1.3% in AP and 1.4% in BWT, underscoring the effectiveness of enhancements to each framework component.
+Notably, even without memory replay, TasLoRA nearly matches the performance of SAPT with memory replay on the Long Sequence Benchmark, particularly in BWT, with only a minor difference of 0.8% (-2.04% vs.
 -1.25%).
-For memory-based methods, both TaSL-M and TasLoRA outperform the best SAPT methods, further demonstrating the versatility and robustness of our framework.
-For the convenience of subsequent experimental analysis and comparison, we focus on the performance of TasLoRA, the best-performing memory-free version, for further evaluations.
+For memory-based methods, both TaSL-M and TasLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework.
+For subsequent experimental analyses and comparisons, we will focus on TasLoRA, the most effective memory-free version, for further evaluations.
 Secondly, TasLoRA consistently demonstrates superior performance across various backbones.
-To further validate the effectiveness of our framework, we conducted experiments using a range of parameter-level backbones, as shown in Figure 4, which highlight performance improvements with increasing model size.
-Across different backbones, our method consistently outperforms traditional approaches.
-For example, in T5-xl, TasLoRA significantly increases the AP metric from 28.6% to 33.3%, and achieves substantial improvements in both FWT and BWT metrics, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
-These results further validate the robust generalization ability of our proposed framework.
+To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones.
+As depicted in Figure 4, performance improvements are evident with increasing model size.
+For instance, in T5-XL, TasLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
+These results further validate the robust generalization ability of our method.
 (a) Results on the SuperNI benchmark
 (b) Results on the Long Sequence benchmark
-Figure 4: Performance of TaSL w/ different backbones.
-Thirdly, adaptive skill consolidation effectively mitigates catastrophic forgetting.
-To rigorously assess our model’s effectiveness in countering forgetting, we analyzed its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark.
-Figure 5 shows that our TasLoRA method significantly slows the forgetting rate, with an average performance decrease of only 11% after training on the final task.
-In stark contrast, vanilla backbones exhibit a substantial average performance drop of 20%, underscoring our method’s superior ability to mitigate forgetting.
-Figure 5: Performance trajectory of Task 1 on Long Sequence Benchmark during the CL process.
+Figure 4: Performance of TasLoRA across various backbones.
+Thirdly, our skill consolidation technique effectively counters catastrophic forgetting.
+To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark.
+Figure 5 illustrates that TasLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task.
+In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
+Figure 5: Performance trajectory of Task 1 on Long Sequence Benchmark throughout the continual learning process.
 VI-C Performance on Unseen Tasks
-Following previous work [16], we further select three tasks from each one of the above task categories to assess the TaSL’s cross-task generalization ability.
-This is also a crucial dimension for evaluating CL algorithms.
-Table III shows the results.
-T5-ZS represents the zero-shot approaches for task adaptation, respectively.
-TasLoRA achieved the best performance, which can be attributed to its ability to identify task-specific and shared parameters effectively.
-TABLE III: Results on unseen tasks based on the T5-Large backbone model. We report the average Rouge-L of the 3 tasks under each category.
+Building on prior research [15], we selected three tasks from each category to evaluate our method’s cross-task generalization capabilities, a critical measure for CL algorithms.
+The results, indicated by the average Rouge-L scores in Table III, highlight the effectiveness of our approach.
+T5-ZS refers to the zero-shot task adaptation methods employed.
+TasLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
+TABLE III: Results on unseen tasks using the T5-Large backbone.
 Unseen Tasks Avg.
@@
 VI-D Visualization of Skill Units
-We visualized the distribution of importance scores for the skill units across tasks on T5-large, as shown in Figure 6 and Figure 7, leading to several critical insights:
-  • There is a noticeable variation in the importance of skill units for the same task, with important skill units making up only a small fraction of all trainable LoRA parameters.
-  • The distribution of important skill units is task-dependent, indicating both task-shared and specific parameters, confirming TaSL’s validity.
-  • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more significant role.
-In contrast, for generative tasks, such as dialogue generation and summarization in the SuperNI benchmark (Figure 7), both the encoder and decoder skill units are important, with both the lower and upper layers of the network being crucial.
-  • Within each layer, the importance of the Query (Q) matrices in the attention mechanism consistently surpasses that of the Value (V) matrices.
-Figure 6: Visualization of importance scores for skill units across different tasks on T5-large for the Long Sequence Benchmark.
-Figure 7: Visualization of importance scores for skill units across different tasks on T5-large for the SuperNI Benchmark.
-VI-E Ablation Study
-In this section, we evaluate the impact of importance-aware skill localization, fine-grained skill consolidation, and hyperparameter sensitivity, with the results discussed below.
-VI-E1 Effect of the proposed importance metric in skill localization
-Our method calculates importance scores using Eq.
-(6).
-As shown in Table IV, we explore alternative importance scoring approaches: (i) to test the effectiveness of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq.
-(6) to include only sensitivity, as described in Eq.
-(9); and (ii) to assess the validity of our proposed approximate second-order gradient importance metric in Eq.
-(6), we use a first-order Taylor expansion to approximate the importance $I$ by using Eq.
-(8).
-The results indicate that using moving averages for importance scoring outperforms the alternatives, with the other two variants leading to performance decreases of up to 3.3%, 2.5%, and 4.2% across the three metrics.
-This underscores the value of accurate skill localization in enhancing model performance.
-TABLE IV: Ablation study. Evaluating the impact of importance scoring variations on skill localization.
+We visualized the distribution of importance scores for skill units across tasks on the T5-large model, as shown in Figure 6 and Figure 7, leading to several critical insights:
+  • There is noticeable variability in the importance of skill units for the same task, with important skill units comprising only a small fraction of all trainable LoRA parameters.
+  • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters.
+This observation substantiates the feasibility of the design motivations behind the TaSL framework.
+  • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more pivotal role.
+Conversely, for generative tasks, like dialogue generation and summarization in the SuperNI benchmark (Figure 7), both encoder and decoder units are crucial, impacting both lower and upper layers of the network.
+  • Within each layer, the importance of the Query (Q) matrices in the attention mechanism consistently exceeds that of the Value (V) matrices.
+Figure 6: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the Long Sequence Benchmark.
+Figure 7: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the SuperNI Benchmark.
+TABLE IV: Ablation Study for Importance-Aware Skill Localization.
 Method AP FWT BWT
@@
 TasLoRA (ours) 72.29 12.89 -2.04
+VI-E Ablation Study
+In this section, we assess the impact of importance-aware skill localization, fine-grained skill consolidation, and memory size on model performance.
+Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
+VI-E1 Effect of the Proposed Importance Metric in Skill Localization
+We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq.
+(6) to focus solely on sensitivity, as outlined in Eq.
+(9); (ii) To verify the validity of our proposed approximate second-order gradient importance metric in Eq.
+(6), we implement a first-order Taylor expansion as described in Eq.
+(8).
+The comparative results, presented in Table IV, show that using moving averages for importance scoring substantially enhances model performance.
+Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics.
+These findings underscore the significant role of accurate skill localization in elevating model effectiveness.
 VI-E2 Effect of Adaptive Skill Consolidation
-We compared our fine-grained skill unit-level adaptive LoRA averaging mechanism against two coarse-grained strategies: (i) Weight-Ensemble, which uses a global weight to uniformly average the LoRA parameters, i.e., Eq.
-(14), and (ii) Exponential Moving Average (EMA) [60], which applies a running average of parameters at each fine-tuning iteration.
-For Table V, Weight-Ensemble significantly improves upon the vanilla model, highlighting the benefits of coarse-grained averaging for continual learning.
-EMA generally outperforms Weight-Ensemble but falls short of our fine-grained approach due to its overuse of averaging, with frequent parameter adjustments within the same task potentially leading to suboptimal outcomes.
-Our method, which averages weights only after each task, enhances computational efficiency.
-TABLE V: Ablation study. Comparing coarse- and fine-grained model averaging methods on skill consolidation.
+We evaluate our fine-grained, skill unit-level adaptive weighted consolidation against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq.
+(14), and (ii) Exponential Moving Average (EMA) [52], which computes a running average of parameters at each fine-tuning iteration.
+According to Table V, Weight-Ensemble significantly enhances the performance of the vanilla model, demonstrating the advantages of coarse-grained averaging in CL.
+While EMA generally outperforms Weight-Ensemble, it does not match the effectiveness of our fine-grained approach.
+EMA’s frequent parameter adjustments within the same task can lead to suboptimal outcomes.
+In contrast, our method, which strategically averages weights only after the completion of each task, boosts computational efficiency and model performance.
+TABLE V: Ablation Study for Fine-Grained Skill Consolidation.
 Method AP FWT BWT
@@
 TasLoRA (ours) 72.29 12.89 -2.04
-VI-E3 Sensitivity Analysis for Hyperparameters
-The proposed framework incorporates three key hyperparameters, including the $\alpha$ for computing importance scores in Eq.
-(10) and Eq.
-(11), the $\beta$ for calculating cumulative importance scores in Eq.
-(13), and the $\tau$ for performing weighted averaging within skill units as outlined in Eq.
-(16).
-Our analysis aims to assess the impact of varying these hyperparameters on our method’s performance, testing on the T5-large backbone model.
-As evidenced in Table VI, we determine that the optimal setting for $\alpha$ is 0.55.
-An $\alpha$ value too low results in a performance decline, indicating that the calculated importance scores are not sufficiently accurate.
-Furthermore, as depicted in Table VII, we also find that $\beta$ values within a normal range do not significantly affect performance.
-However, excessively high or low values for $\beta$ may skew the model towards favoring either past or current task knowledge, thereby disrupting the desired balance.
-Nonetheless, the model’s performance remains relatively stable across most conditions, indicating a low sensitivity to hyperparameter variations.
-The VIII below shows the model’s performance under different temperature coefficients, $\tau$.
-It can be seen that setting a higher $\tau$ (higher than 0.5) smooths the weights, thereby increasing the weight of unimportant skill units, which reduces the model’s effectiveness and may contaminate historical knowledge, leading to forgetting.
-On the other hand, setting a lower $\tau$ (less than 0.1) may cause larger differences in weight coefficients, thereby limiting KT between different tasks.
-TABLE VI: Performance comparisons of TaSL equipped with different $\alpha$ on Long Sequence Benchmark at task order 1.
-$\alpha_{1}$,$\alpha_{2}$ AP FWT BWT
-Vanilla T5-large 51.6 5.6 -29.7
-0.15 71.8 13.7 -3.7
-0.35 71.2 14.1 -4.3
-0.55 72.8 12.6 -2.5
-0.85 70.7 12.9 -3.3
-0.95 71.7 14.0 -3.6
-TABLE VII: Performance comparisons of TaSL equipped with different $\beta$ on Long Sequence Benchmark at task order 1.
-$\beta$ AP FWT BWT
-Vanilla T5-large 51.6 5.6 -29.7
-0.1 73.8 13.6 -3.7
-0.3 73.5 12.4 -3.4
-0.5 74.3 13.5 -2.2
-0.7 72.7 12.9 -2.3
-0.9 60.2 14.2 -5.0
-TABLE VIII: Performance comparisons of TaSL (using T5-large as the backbone) equipped with different $\tau$ on Long Sequence Benchmark at task order 1.
-temperature coefficient $\tau$ AP FWT BWT
-0.05 69.0 14.3 -3.4
-0.15 73.4 12.8 -2.1
-0.25 72.2 13.2 -2.5
-0.35 72.1 13.6 -2.1
-0.45 72.7 12.5 -3.0
-0.55 70.9 11.6 -4.2
+VI-E3 The Effect of Memory Size
+We examine how varying the memory size affects the performance of Replay, TaSL-M, and TasLoRA-M.
+By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI.
+As expected, increasing the memory size generally enhances the performance of all methods, with Replay showing the most significant improvement.
+This is predictable as Replay heavily depends on memory to mitigate catastrophic forgetting.
+When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
+Conversely, TaSL-M and TasLoRA-M utilize skill consolidation strategies to effectively maintain parameters that store historical knowledge.
+Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
+TABLE VI: Ablation Study on Memory Size.
+Memory Size
+2% 5% 10% 50%
+Replay 75.3 76.7 78.4 89.9
+TaSL-M 84.2 85.1 85.7 91.2
+TasLoRA-M 84.3 85.4 86.2 91.8
 VII Conclusion
-In this paper, we introduce a novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning, facilitating effective knowledge transfer across tasks.
-Our framework leverages an importance-aware skill localization and a fine-grained skill consolidation technique to differentiate between task-specific and shared knowledge within each skill unit, thereby mitigating forgetting.
-TaSL also demonstrates strong generalizability and extensibility, with optimizations to various components and the ability to integrate with memory replay methods for further performance enhancement.
-Comprehensive experiments showcase our framework’s exceptional ability to balance preserving past knowledge and excelling in new tasks, surpassing previous state-of-the-art methods.
-While our method significantly improves efficiency and performance in large language model continual learning, some limitations remain.
-For instance, the choice of importance thresholds after determining the importance distribution can affect the effectiveness of subsequent consolidation.
-Dynamically selecting thresholds based on data distribution could lead to more accurate classification of task-shared and task-specific parameters, thereby enhancing performance.
-Additionally, we could explore merging the skill localization and consolidation stages, allowing for parameter consolidation based on importance during model training.
-This would enable more flexible adaptation and address scenarios in online continual learning.
-[Dataset Details] Table IX & X show details of the datasets we used for our experiments, along with their evaluation metrics.
-Overall, in SuperNI, we choose 3 tasks from dialogue generation (Dialog), information extraction (IE), question answering (QA), summarization (Sum) and sentiment analysis (SA), respectively.
-For the Long Sequence benchmark, this includes five tasks from the standard CL benchmark (AG News, Amazon reviews, Yelp reviews, DBpedia and Yahoo Answers), four from GLUE benchmark (MNLI, QQP, RTE, SST2), five from SuperGLUE benchmark (WiC, CB, COPA, MultiRC, BoolQ), and the IMDB movie reviews dataset.
-We report 4 different task orders used for our experiments in Table XI.
-TABLE IX: The details of 15 datasets in the SuperNI Benchmark [48].
-Dataset name Task Metric
-1. task639_multi_woz_user_utterance_generation dialogue generation Rouge-L
-2. task1590_diplomacy_text_generation dialogue generation Rouge-L
-3. task1729_personachat_generate_next dialogue generation Rouge-L
-4. task181_outcome_extraction information extraction Rouge-L
-5. task748_glucose_reverse_cause_event_detection information extraction Rouge-L
-6. task1510_evalution_relation_extraction information extraction Rouge-L
-7. task002_quoref_answer_generation question answering Rouge-L
-8. task073_commonsenseqa_answer_generation question answering Rouge-L
-9. task591_sciq_answer_generation question answering Rouge-L
-10. task511_reddit_tifu_long_text_summarization summarization Rouge-L
-11. task1290_xsum_summarization summarization Rouge-L
-12. task1572_samsum_summary summarization Rouge-L
-13. task363_sst2_polarity_classification sentiment analysis accuracy
-14. task875_emotion_classification sentiment analysis accuracy
-15. task1687_sentiment140_classification sentiment analysis accuracy
-TABLE X: The details of 15 classification datasets in the Long Sequence Benchmark [18]. First five tasks correspond to the standard CL benchmark [61].
-Dataset name Category Task Domain Metric
-1. Yelp CL Benchmark sentiment analysis Yelp reviews accuracy
-2. Amazon CL Benchmark sentiment analysis Amazon reviews accuracy
-3. DBpedia CL Benchmark topic classification Wikipedia accuracy
-4. Yahoo CL Benchmark topic classification Yahoo Q&A accuracy
-5. AG News CL Benchmark topic classification news accuracy
-6. MNLI GLUE natural language inference various accuracy
-7. QQP GLUE paragraph detection Quora accuracy
-8. RTE GLUE natural language inference news, Wikipedia accuracy
-9. SST-2 GLUE sentiment analysis movie reviews accuracy
-10. WiC SuperGLUE word sense disambiguation lexical databases accuracy
-11. CB SuperGLUE natural language inference various accuracy
-12. COPA SuperGLUE question and answering blogs, encyclopedia accuracy
-13. BoolQA SuperGLUE boolean question and answering Wikipedia accuracy
-14. MultiRC SuperGLUE question and answering various accuracy
-15. IMDB SuperGLUE sentiment analysis movie reviews accuracy
-TABLE XI: Four different orders of task sequences used for our experiments. Orders 1-2 correspond to the SuperNI benchmark. Orders 3-4 are long-sequence orders following [18].
-Order Model Task Sequence
-1 T5, LLaMA-2
-task1572 → task363 → task1290 → task181 → task002 →
-task1510 → task639 → task1729 → task073 → task1590 →
-task748 → task511 → task591 → task1687 → task875
-2 T5, LLaMA-2
-task748 → task073 → task1590 → task639 → task1572 →
-task1687 → task591 → task363 → task1510 → task1729 →
-task181 → task511 → task002 → task1290 → task875
-3 T5, LLaMA-2
-mnli → cb → wic → copa → qqp → boolqa → rte → imdb →
-yelp → amazon → sst-2 → dbpedia → ag → multirc → yahoo
-4 T5, LLaMA-2
-yelp → amazon → mnli → cb → copa → qqp → rte → imdb →
-sst-2 → dbpedia → ag → yahoo → multirc → boolqa → wic
+In this paper, we introduced the novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning.
+TaSL employs importance-aware skill localization and fine-grained skill consolidation to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer.
+The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance.
+Extensive testing shows that TaSL excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches.
+Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
+Despite significant improvements in efficiency and performance in large language model continual learning, certain limitations persist.
+For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of skill consolidation.
+Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance.
+Additionally, merging the skill localization and consolidation phases could allow for ongoing parameter consolidation based on their importance during model training.
+This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.

### v3

@@
-TaSL: Task Skill Localization and Consolidation for Language Model Continual Learning
+KlF: Knowledge Localization and Fusion for Language Model Continual Learning
 Abstract
@@
 Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks.
-In this paper, we introduce a novel CL framework for language models, named Task Skill Localization and Consolidation (TaSL), which boosts knowledge transfer without depending on memory replay.
-TaSL initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control.
-Subsequently, it employs a novel group-wise skill localization technique to ascertain the importance distribution of skill units for a new task.
-By comparing this importance distribution with those from previous tasks, we implement a fine-grained skill consolidation strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer.
-As a result, TaSL achieves an optimal balance between retaining prior knowledge and excelling in new tasks.
-TaSL also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA.
+In this paper, we introduce a novel CL framework for language models, named Knowledge Localization and Fusion (KlF), which boosts knowledge transfer without depending on memory replay.
+KlF initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control.
+Subsequently, it employs a novel group-wise knowledge localization technique to ascertain the importance distribution of skill units for a new task.
+By comparing this importance distribution with those from previous tasks, we implement a fine-grained knowledge fusion strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer.
+As a result, KlF achieves an optimal balance between retaining prior knowledge and excelling in new tasks.
+KlF also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA.
 Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques.
-Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of TaSL and its variants across different settings.
+Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of KlF and its variants across different settings.
 Index Terms:
@@
 Achieving a balance between retaining previous knowledge and excelling in new tasks is vital for success.
-Figure 1: Conceptual illustration of TaSL.
+Figure 1: Conceptual illustration of KlF.
 By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
 While these strategies might mitigate CF, they cut off the potential for leveraging knowledge distributed across various PEFT blocks, thus obstructing bi-directional KT among tasks and resulting in less-than-optimal performance (Limitation 2).
-To address these limitations, we introduce Task Skill Localization and Consolidation (TaSL) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay.
+To address these limitations, we introduce Knowledge Localization and Fusion (KlF) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay.
 Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20].
@@
 Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
-Based on these insights, TaSL facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks.
-Initially, TaSL utilizes a group-wise importance-aware skill localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
+Based on these insights, KlF facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks.
+Initially, KlF utilizes a group-wise importance-aware knowledge localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
 By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1.
-The subsequent skill consolidation phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
-In detail, TaSL first reconstructs the model or PEFT block into fine-grained “skill units”.
+The subsequent knowledge fusion phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
+In detail, KlF first reconstructs the model or PEFT block into fine-grained “skill units”.
 A skill unit refers to a distinct subset of model parameters that encapsulates specific functional capabilities or knowledge relevant to a particular task, such as the Query matrix within the self-attention layer.
 By operating at this finer granularity, we can localize and consolidate task-specific and shared knowledge within a single PEFT block, rather than adapting a separate PEFT block for each task as in previous works (addressing Limitation 1).
-The importance-aware skill localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task.
-Our skill consolidation phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge.
+The importance-aware knowledge localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task.
+Our knowledge fusion phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge.
 This stage facilitates forward KT by using previously fine-tuned weights as the starting point for new tasks.
@@
 To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
-Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the TaSL framework specifically for LoRA.
-While the initial TaSL framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored.
-Additionally, the reliance on first-order gradients for skill localization in TaSL may not yield precise importance localization.
-The skill consolidation phase also involves many hyperparameters, which could complicate the averaging process.
-To address these issues, we propose TasLoRA, a variant of TaSL optimized for LoRA.
-TasLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning.
+Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the KlF framework specifically for LoRA.
+While the initial KlF framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored.
+Additionally, the reliance on first-order gradients for knowledge localization in KlF may not yield precise importance localization.
+The knowledge fusion phase also involves many hyperparameters, which could complicate the averaging process.
+To address these issues, we propose KlFLoRA, a variant of KlF optimized for LoRA.
+KlFLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning.
 It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter.
-For skill localization, TasLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores.
-In skill consolidation, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit.
-Additionally, we combine TaSL with memory replay to broaden its application scope, introducing the TaSL-M model.
-Through extensive experiments on two CL benchmarks with four parameter-level backbones, our TaSL framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
+For knowledge localization, KlFLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores.
+In knowledge fusion, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit.
+Additionally, we combine KlF with memory replay to broaden its application scope, introducing the KlF-M model.
+KlF framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
 The main contributions are summarized as follows:
-  • We introduce the Task Skill Localization and Consolidation (TaSL) framework for language model continual learning.
-By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, TaSL ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
-  • We devise various group-wise skill localization and fine-grained skill consolidation techniques.
-Notably, our skill localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance.
-For skill consolidation, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
-  • The TaSL framework is distinguished by its robust generalizability and extensibility.
+  • We introduce the Knowledge Localization and Fusion (KlF) framework for language model continual learning.
+By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, KlF ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
+  • We devise various group-wise knowledge localization and fine-grained knowledge fusion techniques.
+Notably, our KlF localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance.
+For knowledge fusion, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
+  • The KlF framework is distinguished by its robust generalizability and extensibility.
 Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures.
-Moreover, integrating TaSL with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
-  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our TaSL framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
-Furthermore, TaSL consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
+Moreover, integrating KlF with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
+  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our KlF framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
+Furthermore, KlF consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
 II Related Work
@@
 While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks.
-Our TaSL framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting.
-To further enhance performance, we extend TaSL with memory replay techniques, resulting in TaSL-M, improving its flexibility and applicability across various scenarios.
+Our KlF framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting.
+To further enhance performance, we extend KlF with memory replay techniques, resulting in KlF-M, improving its flexibility and applicability across various scenarios.
 Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy.
@@
 This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting.
-To address these challenges in the context of PEFT-based CL, we extend the TaSL framework and introduce TasLoRA.
-By dividing fine-grained skill units based on parameter dependencies within LoRA, TasLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
-II-B Skill Localization
+To address these challenges in the context of PEFT-based CL, we extend the KlF framework and introduce KlFLoRA.
+By dividing fine-grained skill units based on parameter dependencies within LoRA, KlFLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
+II-B Knowledge Localization
 Research has shown that model parameters contribute unevenly to performance [32].
@@
 While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency.
-In contrast, our importance-aware skill localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time.
-Combined with our skill consolidation strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
+In contrast, our importance-aware knowledge localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time.
+Combined with our knowledge fusion strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
 III Preliminaries
@@
 And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging.
-Our TaSL framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18].
-To broaden its applicability, we have augmented TaSL with memory replay.
-In this enhanced setup, known as TaSL-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$.
+Our KlF framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18].
+To broaden its applicability, we have augmented KlF with memory replay.
+In this enhanced setup, known as KlF-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$.
 The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.
@@
 Due to its effectiveness, LoRA has emerged as one of the most favored PEFT methods within the NLP community [35].
-IV Proposed Method: TaSL
+IV Proposed Method: KlF
 IV-A Overview
-TaSL includes two key components: (i) Skill Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Skill Consolidation, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
-Figure 2 presents a detailed overview of our proposed TaSL framework, with the following subsections elaborating on each component and the corresponding extensions.
+KlF includes two key components: (i) Knowledge Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Knowledge Fusion, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
+Figure 2 presents a detailed overview of our proposed KlF framework, with the following subsections elaborating on each component and the corresponding extensions.
 Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units.
-Right: Overview of TaSL.
-Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware skill localization method during fine-tuning.
-Step 2: Based on a categorical model averaging strategy, the skill consolidation mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$.
+Right: Overview of KlF.
+Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware knowledge localization method during fine-tuning.
+Step 2: Based on a categorical model averaging strategy, the knowledge fusion mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$.
 This integration is strategically guided by the importance distributions of skill units across different tasks.
@@
 In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA.
-This approach aligns with the original TaSL framework and offers the advantage of being model-agnostic.
+This approach aligns with the original KlF framework and offers the advantage of being model-agnostic.
 It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
@@
 where $I$ is the identity matrix, this forces $A$ and $B$ to be orthogonal after training.
-IV-C Importance-aware Skill Localization
+IV-C Importance-aware Knowledge Localization
 To calculate the importance of each skill unit $u$, we introduce a group-wise metric that mitigates the significant computational and storage burdens associated with previous parameter-level importance calculation methods [37]:
@@
 IV-C1 First-Order Gradient-Based Metric
-In the original TaSL framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
+In the original KlF framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
 $I_{W_{i}}=\left|W_{i}\nabla_{W_{i}}\mathcal{L}\right|$ (8)
@@
 The detailed computation process is provided in Algorithm 1.
-Algorithm 1 Importance-aware Skill Localization
+Algorithm 1 Importance-aware Knowledge Localization
 0:  Training dataset $\mathcal{D}_{k}$ for task $\mathcal{T}_{k}$; total training iterations $T$; hyperparameters $\alpha_{1},\alpha_{2}$.
@@
 The initial scores, $\mathcal{I}(\hat{\mathcal{U}}_{1})$, are set equal to $\mathcal{I}(\mathcal{U}_{1})$.
-Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the skill consolidation process.
-IV-D Skill Consolidation
-Following skill localization, the next critical phase is skill consolidation, where knowledge within each skill unit is integrated into a cohesive framework.
+Following this, the importance distribution for skill units up to task $\mathcal{T}_{k-1}$ is combined with that of the current task, $\mathcal{T}_{k}$, to facilitate the knowledge fusion process.
+IV-D Knowledge Fusion
+Following knowledge localization, the next critical phase is knowledge fusion, where knowledge within each skill unit is integrated into a cohesive framework.
 This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
@@
 This refined approach offers a tailored solution, addressing the unique requirements of each task more effectively.
-IV-D1 Static Weighted Consolidation
-In the original TaSL framework, this consolidation strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores.
+IV-D1 Static Weighted Fusion
+In the original KlF framework, this merging strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores.
 A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
-The static weighted consolidation strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows:
+The static weighted fusion strategy tailors parameter integration for each skill unit, according to its importance across different tasks. The method is defined as follows:
 $\hat{u}_{i}^{k}=\begin{cases}\gamma\hat{u}_{i}^{k-1}+(1-\gamma)u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{+}$}\\ \hat{u}_{i}^{k-1},&\text{if $(\hat{u}_{i}^{k-1})^{+}$, $(u_{i}^{k})^{-}$}\\ u_{i}^{k},&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{+}$}\\ \frac{1}{2}(\hat{u}_{i}^{k-1}+u_{i}^{k}),&\text{if $(\hat{u}_{i}^{k-1})^{-}$, $(u_{i}^{k})^{-}$}\\ \end{cases}$ (15)
@@
 The hyperparameter $\gamma$ is employed to control their influences.
-IV-D2 Adaptive Weighted Consolidation
+IV-D2 Adaptive Weighted Fusion
 As a sophisticated extension to the static method, which necessitated the manual setting of multiple hyperparameters, this adaptive strategy simplifies the process by automatically adjusting to various scenarios.
@@
 $\begin{split}\hat{u}_{i}^{k}=\left(\frac{\hat{\lambda}^{k-1}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot\hat{u}_{i}^{k-1}+\left(\frac{\lambda^{k}_{i}}{\hat{\lambda}^{k-1}_{i}+\lambda^{k}_{i}}\right)\cdot u_{i}^{k}\end{split}$ (17)
-By applying Eq. (15) or Eq. (17) in our skill consolidation phrase, it effectively addresses the challenges in CL:
+By applying Eq. (15) or Eq. (17) in our knowledge fusion phrase, it effectively addresses the challenges in CL:
   • Mitigating CF: When $\mathcal{I}(\hat{u}^{k-1}_{i})$ is high and $\mathcal{I}(u^{k}_{i})$ is low, this indicates that the skill unit $u_{i}$ is more critical for previous tasks.
@@
   • Handling Low Relevance: If both $\mathcal{I}(\hat{u}^{k-1}{i})$ and $\mathcal{I}(u^{k}{i})$ are low, indicating minimal relevance to any task, a simple averaging of parameters within this unit is adequate.
-Skill consolidation is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization.
+Knowledge fusion is conducted before starting a new task in CL, utilizing the averaged model for subsequent task initialization.
 Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$.
-Detailed implementation of entire TaSL algorithm is provided in Algorithm 2.
-Algorithm 2 TaSL
+Detailed implementation of entire KlF algorithm is provided in Algorithm 2.
+Algorithm 2 KlF
 0:  Dataset $\mathcal{D}_{k}$ for task $k$ = $1,\ldots,K$; initial pre-trained model $f_{0}$; hyperparameters $\alpha_{1},\alpha_{2},\beta,\tau$.
@@
 2:  for task $k$ = $1,\ldots,K$ do
-3:     # skill localization.
+3:     # knowledge localization.
 4:     Get $f_{k}$ and calculate $\mathcal{U}_{k}$ by Algorithm (1);
@@
 8:     else
-9:        # skill consolidation.
+9:        # knowledge fusion.
 10:        for skill unit $i$ = $1,\ldots,n$ do
@@
 16:  end for
-V TaSL with Memory Replay: TaSL-M
-To adapt TaSL for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, TaSL-M, designed to further improve model performance.
+V KlF with Memory Replay: KlF-M
+To adapt KlF for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, KlF-M, designed to further improve model performance.
 The implementation process is shown in Figure 3.
-Figure 3: Overview of TaSL-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of TaSL.
-Following the training on task $\mathcal{T}_{k}$ using the skill localization and skill consolidation, we obtain the model $\hat{f}_{k}$.
+Figure 3: Overview of KlF-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of KlF.
+Following the training on task $\mathcal{T}_{k}$ using the knowledge localization and fusion, we obtain the model $\hat{f}_{k}$.
 Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$.
@@
 The experiments include two different task orders for each benchmark.
-Further details about the tasks and their order are provided in the Appendix.
 TABLE I: The overall results on two continual learning benchmarks with T5-large model. All results are averaged over two different orders of each benchmark.
@@
 O-LoRA [18] 25.89 8.14 -24.59 69.24 10.15 -4.05
-TaSL (ours) 26.41 11.78 -18.55 70.71 11.80 -3.27
-TasLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -2.04
-Replay 35.37 2.35 -15.79 75.28 3.28 -1.88
+KlF (ours) 26.41 11.78 -18.55 70.71 11.80 -3.67
+KlFLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -3.04
+Replay 35.37 2.35 -15.79 70.98 3.28 -15.42
 SAPT [15] 51.54 8.88 -0.57 82.02 9.86 -1.25
-TaSL-M (ours) 52.12 11.13 -1.31 84.26 12.32 -1.08
-TasLoRA-M (ours) 53.01 11.21 -0.81 84.33 11.71 -0.98
+KlF-M (ours) 52.12 11.13 -1.31 84.21 12.32 -1.08
+KlFLoRA-M (ours) 53.01 11.21 -0.81 84.36 11.71 -0.98
 VI-A2 Evaluation Metrics
@@
 VI-A3 Baselines
-We evaluate TaSL against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
+We evaluate KlF against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
 IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks.
@@
 SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
-Table II compares our TaSL with these baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities.
-To clarify the specific localization and consolidation techniques used in TaSL and its variants: TaSL: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for skill localization (Eq.
-8), and employs static weighted consolidation (Eq.
+Table II compares our KlF with other baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities.
+To clarify the specific localization and fusion techniques used in KlF and its variants: KlF: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for knowledge localization (Eq.
+8), and employs static weighted fusion (Eq.
 15).
-TasLoRA: as the most recent extension of TaSL, this variant introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric for skill localization (Eq.
-9), and features adaptive weighted consolidation (Eq.
+KlFLoRA: an extension of KlF that introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric (Eq.
+9), and features adaptive weighted fusion (Eq.
 17).
-TasL-M and TasLoRA-M: these enhancements of TaSL and TasLoRA integrate memory replay to further boost performance.
-TABLE II: The comparison between TaSL and other CL methods.
+KlF-M and KlFLoRA-M: these enhancements of KlF and KlFLoRA integrate memory replay to further boost performance.
+TABLE II: The comparison between KlF and other CL methods.
 Specifically, RF indicates whether the method is rehearsal-free.
@@
 SAPT [15]
-TaSL
+KlF
 VI-A4 Implementation Details
-We utilize two distinct language model architectures previously adopted in the field of CL for NLP: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50].
-In TaSL, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq.
+We utilize two distinct language model architectures: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50].
+In KlF, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq.
 (10) and Eq.
@@
 Following [51], we allocate 2% of the original training set volume for replay samples in all replay-based baseline methods.
-For different backbones, we utilized the following hyperparameters:
+Hyperparameters are as follows:
   • T5-base (220M), T5-large (780M), and T5-XL (3B): Training was conducted with a learning rate of 3e-4, batch size of 8, maximum input length of 512, maximum target length of 128, a rank of 8, targeting modules [q, v] and 10 epochs.
@@
 For testing, settings included temperature = 0.02, top_p = 0, top_k = 1, max new tokens = 128.
-Experiments are conducted using 4 NVIDIA A100 GPUs.
 VI-B Main Results
 The overall CL results of various methods using the same T5-large backbone are summarized in Table I.
-Firstly, TaSL and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer.
+Firstly, KlF and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer.
 Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
-Furthermore, our enhanced version, TasLoRA, significantly outperforms TaSL.
-Specifically, TasLoRA achieves average gains of 1.3% in AP and 1.4% in BWT, underscoring the effectiveness of enhancements to each framework component.
-Notably, even without memory replay, TasLoRA nearly matches the performance of SAPT with memory replay on the Long Sequence Benchmark, particularly in BWT, with only a minor difference of 0.8% (-2.04% vs.
--1.25%).
-For memory-based methods, both TaSL-M and TasLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework.
-For subsequent experimental analyses and comparisons, we will focus on TasLoRA, the most effective memory-free version, for further evaluations.
-Secondly, TasLoRA consistently demonstrates superior performance across various backbones.
+Furthermore, our enhanced version, KlFLoRA, significantly outperforms KlF.
+Specifically, KlF-M achieves average gains of 1.3% in AP and 0.3% in BWT, underscoring the effectiveness of enhancements to each framework component.
+Notably, even without memory replay, KlF-M surpasses the performance of vanilla repla on the Long Sequence Benchmark, particularly in BWT.
+For the memory-based methods, to ensure a fair comparison, we adopted the same data synthesis-based generative replay approach used in SAPT, ensuring that the amount of data used in memory replay is consistent across methods.
+The results show that both KlF-M and KlFLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework.
+For subsequent experimental analyses and comparisons, we will focus on KlFLoRA, the most effective memory-free version, for further evaluations.
+Secondly, KlFLoRA consistently demonstrates superior performance across various backbones.
 To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones.
 As depicted in Figure 4, performance improvements are evident with increasing model size.
-For instance, in T5-XL, TasLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
+For instance, in T5-XL, KlFLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
 These results further validate the robust generalization ability of our method.
@@
 (b) Results on the Long Sequence benchmark
-Figure 4: Performance of TasLoRA across various backbones.
-Thirdly, our skill consolidation technique effectively counters catastrophic forgetting.
+Figure 4: Performance of KlFLoRA across various backbones.
+Thirdly, our knowledge fusion technique effectively counters catastrophic forgetting.
 To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark.
-Figure 5 illustrates that TasLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task.
+Figure 5 illustrates that KlFLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task.
 In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
@@
 T5-ZS refers to the zero-shot task adaptation methods employed.
-TasLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
+KlFLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
 TABLE III: Results on unseen tasks using the T5-Large backbone.
@@
 LFPT5 6.96 35.32 35.00 13.26 21.51 22.41
-TasLoRA 10.32 31.34 37.13 14.20 47.17 28.03
+KlFLoRA 10.32 31.34 37.13 14.20 47.17 28.03
 VI-D Visualization of Skill Units
@@
   • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters.
-This observation substantiates the feasibility of the design motivations behind the TaSL framework.
+This observation substantiates the feasibility of the design motivations behind the KlF framework.
   • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more pivotal role.
@@
 Figure 7: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the SuperNI Benchmark.
-TABLE IV: Ablation Study for Importance-Aware Skill Localization.
+TABLE IV: Ablation Study for Importance-Aware Knowledge Localization.
 Method AP FWT BWT
@@
 $s\left(\cdot\right)=\left|\nabla_{w_{ij}}\mathcal{L}\right|$ 68.82 10.80 -6.22
-TasLoRA (ours) 72.29 12.89 -2.04
+KlFLoRA (ours) 72.29 12.89 -2.04
 VI-E Ablation Study
-In this section, we assess the impact of importance-aware skill localization, fine-grained skill consolidation, and memory size on model performance.
+In this section, we assess the impact of importance-aware knowledge localization, fine-grained knowledge fusion, and memory size on model performance.
 Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
-VI-E1 Effect of the Proposed Importance Metric in Skill Localization
+VI-E1 Effect of the Proposed Importance Metric in Knowledge Localization
 We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq.
@@
 Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics.
-These findings underscore the significant role of accurate skill localization in elevating model effectiveness.
-VI-E2 Effect of Adaptive Skill Consolidation
-We evaluate our fine-grained, skill unit-level adaptive weighted consolidation against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq.
+These findings underscore the significant role of accurate knowledge localization.
+VI-E2 Effect of Adaptive Knowledge Fusion
+We evaluate our fine-grained, skill unit-level adaptive weighted fusion against two coarse-grained strategies: (i) Weight-Ensemble, which uniformly averages the LoRA parameters using a global weight as per Eq.
 (14), and (ii) Exponential Moving Average (EMA) [52], which computes a running average of parameters at each fine-tuning iteration.
@@
 In contrast, our method, which strategically averages weights only after the completion of each task, boosts computational efficiency and model performance.
-TABLE V: Ablation Study for Fine-Grained Skill Consolidation.
+TABLE V: Ablation Study for Fine-Grained Knowledge Fusion.
 Method AP FWT BWT
@@
 EMA 62.76 8.23 -9.80
-TasLoRA (ours) 72.29 12.89 -2.04
+KlFLoRA (ours) 72.29 12.89 -3.04
 VI-E3 The Effect of Memory Size
-We examine how varying the memory size affects the performance of Replay, TaSL-M, and TasLoRA-M.
+We examine how varying the memory size affects the performance of Replay, KlF-M, and KlFLoRA-M.
 By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI.
@@
 When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
-Conversely, TaSL-M and TasLoRA-M utilize skill consolidation strategies to effectively maintain parameters that store historical knowledge.
+Conversely, KlF-M and KlFLoRA-M utilize knowledge fusion strategies to effectively maintain parameters that store historical knowledge.
 Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
@@
 2% 5% 10% 50%
-Replay 75.3 76.7 78.4 89.9
-TaSL-M 84.2 85.1 85.7 91.2
-TasLoRA-M 84.3 85.4 86.2 91.8
+Replay 71.2 72.4 73.8 76.1
+KlF-M 74.2 75.1 75.7 79.2
+KlFLoRA-M 74.3 75.4 76.2 79.8
 VII Conclusion
-In this paper, we introduced the novel Task Skill Localization and Consolidation (TaSL) framework for language model continual learning.
-TaSL employs importance-aware skill localization and fine-grained skill consolidation to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer.
+In this paper, we introduced the novel Knowledge Localization and Fusion (KlF) framework for language model continual learning.
+KlFLoRA employs importance-aware knowledge localization and fine-grained knowledge fusion to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer.
 The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance.
-Extensive testing shows that TaSL excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches.
+Extensive testing shows that KlFLoRA excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches.
 Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
 Despite significant improvements in efficiency and performance in large language model continual learning, certain limitations persist.
-For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of skill consolidation.
+For instance, the choice of importance thresholds following the determination of importance distribution can impact the effectiveness of knowledge fusion.
 Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance.
-Additionally, merging the skill localization and consolidation phases could allow for ongoing parameter consolidation based on their importance during model training.
+Additionally, merging the knowledge localization and fusion phases could allow for ongoing parameter consolidation based on their importance during model training.
 This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.

### v4

@@
-KlF: Knowledge Localization and Fusion for Language Model Continual Learning
+KIF: Knowledge Identification and Fusion for Language Model Continual Learning
 Abstract
@@
 Existing approaches commonly utilize multiple parameter-efficient fine-tuning (PEFT) blocks to acquire task-specific knowledge, yet these methods are inefficient and fail to leverage potential knowledge transfer across tasks.
-In this paper, we introduce a novel CL framework for language models, named Knowledge Localization and Fusion (KlF), which boosts knowledge transfer without depending on memory replay.
-KlF initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control.
-Subsequently, it employs a novel group-wise knowledge localization technique to ascertain the importance distribution of skill units for a new task.
+In this paper, we introduce a novel CL framework for language models, named Knowledge Iidentification and Fusion (KIF), which boosts knowledge transfer without depending on memory replay.
+KIF initially segregates the model into ‘skill units’ based on parameter dependencies, allowing for more precise control.
+Subsequently, it employs a novel group-wise knowledge identification technique to ascertain the importance distribution of skill units for a new task.
 By comparing this importance distribution with those from previous tasks, we implement a fine-grained knowledge fusion strategy that retains task-specific knowledge, thereby preventing forgetting, and updates task-shared knowledge, which facilitates bi-directional knowledge transfer.
-As a result, KlF achieves an optimal balance between retaining prior knowledge and excelling in new tasks.
-KlF also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA.
+As a result, KIF achieves an optimal balance between retaining prior knowledge and excelling in new tasks.
+KIF also demonstrates strong generalizability, making it suitable for various base models and adaptable to PEFT methods like LoRA.
 Furthermore, it offers notable extensibility, supporting enhancements through integration with memory replay techniques.
-Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of KlF and its variants across different settings.
+Comprehensive experiments conducted on two CL benchmarks, involving models ranging from 220M to 7B parameters, affirm the effectiveness of KIF and its variants across different settings.
 Index Terms:
@@
 Achieving a balance between retaining previous knowledge and excelling in new tasks is vital for success.
-Figure 1: Conceptual illustration of KlF.
+Figure 1: Conceptual illustration of KIF.
 By identifying task-relevant areas across both previously accumulated and current tasks, we can consolidate the task-shared and task-specific parameters to facilitate efficient knowledge transfer (KT) and mitigate catastrophic forgetting (CF).
@@
 While these strategies might mitigate CF, they cut off the potential for leveraging knowledge distributed across various PEFT blocks, thus obstructing bi-directional KT among tasks and resulting in less-than-optimal performance (Limitation 2).
-To address these limitations, we introduce Knowledge Localization and Fusion (KlF) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay.
+To address these limitations, we introduce Knowledge Identification and Fusion (KIF) [19], a novel CL framework designed to improve KT between tasks without relying on memory replay.
 Our approach is motivated by recent findings that model parameters contribute unevenly to performance [20].
@@
 Additionally, findings from [14] indicate that fine-tuning with LoRA often preserves many unnecessary parameter changes, pointing to substantial redundancies within PEFT blocks.
-Based on these insights, KlF facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks.
-Initially, KlF utilizes a group-wise importance-aware knowledge localization technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
+Based on these insights, KIF facilitates KT by pinpointing and consolidating the importance distribution of model parameters across tasks.
+Initially, KIF utilizes a group-wise importance-aware knowledge identification technique that employs gradient trajectories to identify tiny regions within the parameter space that store crucial knowledge for the current task.
 By comparing the importance distribution with those of previous tasks, we can then differentiate between task-specific and task-shared regions, as illustrated in Figure 1.
 The subsequent knowledge fusion phase then categorically integrates weights from previous tasks with the current one, efficiently mitigating CF.
-In detail, KlF first reconstructs the model or PEFT block into fine-grained “skill units”.
+In detail, KIF first reconstructs the model or PEFT block into fine-grained “skill units”.
 A skill unit refers to a distinct subset of model parameters that encapsulates specific functional capabilities or knowledge relevant to a particular task, such as the Query matrix within the self-attention layer.
 By operating at this finer granularity, we can localize and consolidate task-specific and shared knowledge within a single PEFT block, rather than adapting a separate PEFT block for each task as in previous works (addressing Limitation 1).
-The importance-aware knowledge localization method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task.
+The importance-aware knowledge identification method employs a new group-wise metric to compute importance scores, effectively quantifying the significance of each skill unit for the current task.
 Our knowledge fusion phase, then based on a fine-grained model averaging strategy, effectively manages different types of knowledge.
@@
 To prevent CF, we ensure the integrity of skill units containing prior task-specific knowledge is maintained, safeguarding them from being altered by the learning of new tasks (addressing Limitation 2).
-Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the KlF framework specifically for LoRA.
-While the initial KlF framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored.
-Additionally, the reliance on first-order gradients for knowledge localization in KlF may not yield precise importance localization.
+Given the widespread adoption and success of LoRA in fine-tuning LLMs, there is a compelling opportunity to optimize the KIF framework specifically for LoRA.
+While the initial KIF framework has shown promising results, its validation has been limited to a specific sub-task, and its broader applicability on general CL benchmarks remains explored.
+Additionally, the reliance on first-order gradients for knowledge identification in KIF may not yield precise importance localization.
 The knowledge fusion phase also involves many hyperparameters, which could complicate the averaging process.
-To address these issues, we propose KlFLoRA, a variant of KlF optimized for LoRA.
-KlFLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning.
+To address these issues, we propose KIFLoRA, a variant of KIF optimized for LoRA.
+KIFLoRA redesigns the LoRA adapter into new skill units based on parameter dependencies, streamlining knowledge management throughout sequential task learning.
 It incorporates an orthogonal loss during fine-tuning to ensure distinct latent semantic features are captured within the same LoRA adapter.
-For knowledge localization, KlFLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores.
+For knowledge identification, KIFLoRA utilizes a new second-order gradient approximate group-wise metric, enhancing the precision of importance scores.
 In knowledge fusion, it shifts from hard-mask to soft-mask averaging, employing an adaptive technique that flexibly integrates task-specific and shared parameters according to the importance of the skill unit.
-Additionally, we combine KlF with memory replay to broaden its application scope, introducing the KlF-M model.
-KlF framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
+Additionally, we combine KIF with memory replay to broaden its application scope, introducing the KIF-M model.
+KIF framework and its variants excel in mitigating CF and showcase remarkable capabilities for KT, outperforming SOTA methods.
 The main contributions are summarized as follows:
-  • We introduce the Knowledge Localization and Fusion (KlF) framework for language model continual learning.
-By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, KlF ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
-  • We devise various group-wise knowledge localization and fine-grained knowledge fusion techniques.
-Notably, our KlF localization utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance.
+  • We introduce the Knowledge Identification and Fusion (KIF) framework for language model continual learning.
+By pinpointing and amalgamating task-specific and task-shared knowledge at a granular skill unit level, KIF ensures effective knowledge transfer and significantly reduces catastrophic forgetting, addressing the shortcomings of previous methodologies.
+  • We devise various group-wise knowledge identification and fine-grained knowledge fusion techniques.
+Notably, our KIF utilizes either first-order gradients or an innovative second-order gradient approximation for assessing parameter importance.
 For knowledge fusion, we implement strategies ranging from categorical hard-masking to adaptive soft-masking, enhancing the flexibility and precision of knowledge integration.
-  • The KlF framework is distinguished by its robust generalizability and extensibility.
+  • The KIF framework is distinguished by its robust generalizability and extensibility.
 Its skill units are designed flexibly, allowing for seamless tailoring to PEFT methods, such as LoRA, optimizing performance across different model architectures.
-Moreover, integrating KlF with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
-  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our KlF framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
-Furthermore, KlF consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
+Moreover, integrating KIF with memory replay enables further performance enhancements and broadens its applicability to diverse scenarios.
+  • Extensive evaluation on two CL benchmarks demonstrates the superiority of our KIF framework and its variants in facilitating knowledge transfer and mitigating catastrophic forgetting, especially in memory-free scenarios.
+Furthermore, KIF consistently excels across diverse model sizes (from 220M to 7B), various model architectures (T5 and LLaMA-2), and unseen tasks.
 II Related Work
@@
 While effective at minimizing forgetting, these methods often hinder knowledge transfer (KT) between tasks.
-Our KlF framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting.
-To further enhance performance, we extend KlF with memory replay techniques, resulting in KlF-M, improving its flexibility and applicability across various scenarios.
+Our KIF framework stands out by bi-directionally analyzing parameter importance across both historical and current tasks, enabling effective KT while mitigating forgetting.
+To further enhance performance, we extend KIF with memory replay techniques, resulting in KIF-M, improving its flexibility and applicability across various scenarios.
 Language model continual learning with PEFT methods in the era of LLMs often adopt a parameter isolation strategy.
@@
 This can lead to overemphasizing unimportant weights, contaminating previously acquired task-specific knowledge, and causing forgetting.
-To address these challenges in the context of PEFT-based CL, we extend the KlF framework and introduce KlFLoRA.
-By dividing fine-grained skill units based on parameter dependencies within LoRA, KlFLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
-II-B Knowledge Localization
+To address these challenges in the context of PEFT-based CL, we extend the KIF framework and introduce KIFLoRA.
+By dividing fine-grained skill units based on parameter dependencies within LoRA, KIFLoRA enhances KT efficiency, enables more precise control, and effectively mitigates forgetting.
+II-B Knowledge Identification
 Research has shown that model parameters contribute unevenly to performance [32].
@@
 While [20] introduced the concept of “skill localization” to identify critical parameters in pre-trained language models, their approach requires additional steps to identify and retrain these parameters post-fine-tuning, which impacts efficiency.
-In contrast, our importance-aware knowledge localization method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time.
+In contrast, our importance-aware knowledge identification method leverages trajectory gradients during each task’s training phase to identify parameter importance, significantly reducing processing time.
 Combined with our knowledge fusion strategy, which explicitly targets task-specific and shared parameters, our approach effectively mitigates forgetting and achieves knowledge transfer in CL.
@@
 And we denote $f_{k}$ as the model trained up to task $\mathcal{T}_{k}$, while $\hat{f}_{k}$ represents the model after integrating knowledge from $\hat{f}_{k-1}$ and $f_{k}$ through averaging.
-Our KlF framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18].
-To broaden its applicability, we have augmented KlF with memory replay.
-In this enhanced setup, known as KlF-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$.
+Our KIF framework is designed to tackle a more challenging scenario where the model cannot access any historical data during training [18].
+To broaden its applicability, we have augmented KIF with memory replay.
+In this enhanced setup, known as KIF-M, we store a random subset of $\left|\mathcal{M}\right|$ samples from each prior task’s training set $\mathcal{D}_{i}$ in memory $\mathcal{M}_{i}$.
 The model is then trained jointly on the new task data $\mathcal{D}_{k}$ and the accumulated memory $\mathcal{M}_{<k}$, leveraging this memory replay to improve performance.
@@
 Due to its effectiveness, LoRA has emerged as one of the most favored PEFT methods within the NLP community [35].
-IV Proposed Method: KlF
+IV Proposed Method: KIF
 IV-A Overview
-KlF includes two key components: (i) Knowledge Localization, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Knowledge Fusion, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
-Figure 2 presents a detailed overview of our proposed KlF framework, with the following subsections elaborating on each component and the corresponding extensions.
+KIF includes two key components: (i) Knowledge Identification, utilizing a group-wise importance metric to precisely identify the importance distribution of parameters across tasks, and (ii) Knowledge Fusion, which employs an innovative fine-grained model averaging strategy to integrate model weights from both current and previous tasks for effective knowledge transfer.
+Figure 2 presents a detailed overview of our proposed KIF framework, with the following subsections elaborating on each component and the corresponding extensions.
 Figure 2: Left: Depiction of reconstructing LoRA as fine-grained skill units.
-Right: Overview of KlF.
-Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware knowledge localization method during fine-tuning.
+Right: Overview of KIF.
+Step 1: We compute the importance scores of skill units for the current task $k$ using our importance-aware knowledge identification method during fine-tuning.
 Step 2: Based on a categorical model averaging strategy, the knowledge fusion mechanism merges the model $\hat{f}_{k-1}$, which accumulates knowledge of all previous tasks, with the current task’s model $f_{k}$.
@@
 In this division strategy, we define skill units as individual matrices in the model, such as the query or key matrices in the self-attention layer or the A matrix in LoRA.
-This approach aligns with the original KlF framework and offers the advantage of being model-agnostic.
+This approach aligns with the original KIF framework and offers the advantage of being model-agnostic.
 It is suitable for both traditional full-parameter fine-tuning and modern parameter-efficient fine-tuning methods like LoRA, ensuring broad generalizability.
@@
 where $I$ is the identity matrix, this forces $A$ and $B$ to be orthogonal after training.
-IV-C Importance-aware Knowledge Localization
+IV-C Importance-aware Knowledge Identification
 To calculate the importance of each skill unit $u$, we introduce a group-wise metric that mitigates the significant computational and storage burdens associated with previous parameter-level importance calculation methods [37]:
@@
 IV-C1 First-Order Gradient-Based Metric
-In the original KlF framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
+In the original KIF framework, we defined the importance function $s(\cdot)$ as the magnitude of the gradient-weight product:
 $I_{W_{i}}=\left|W_{i}\nabla_{W_{i}}\mathcal{L}\right|$ (8)
@@
 However, the significance of a parameter is also influenced by second-order gradients, represented by the Hessian matrix.
-Overlooking this second-order gradient information can introduce biases in the localization process, potentially degrading model performance.
+Overlooking this second-order gradient information can introduce biases in the identification process, potentially degrading model performance.
 Therefore, we propose the following new metric.
@@
 The detailed computation process is provided in Algorithm 1.
-Algorithm 1 Importance-aware Knowledge Localization
+Algorithm 1 Importance-aware Knowledge Identification
 0:  Training dataset $\mathcal{D}_{k}$ for task $\mathcal{T}_{k}$; total training iterations $T$; hyperparameters $\alpha_{1},\alpha_{2}$.
@@
 IV-D Knowledge Fusion
-Following knowledge localization, the next critical phase is knowledge fusion, where knowledge within each skill unit is integrated into a cohesive framework.
+Following knowledge identification, the next critical phase is knowledge fusion, where knowledge within each skill unit is integrated into a cohesive framework.
 This process utilizes a sophisticated model averaging approach that accounts for various factors to optimize task performance.
@@
 IV-D1 Static Weighted Fusion
-In the original KlF framework, this merging strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores.
+In the original KIF framework, this merging strategy initiates by establishing importance thresholds $\delta$ through quantiles, selecting the top 20% of skill units based on importance scores.
 A skill unit $u_{i}^{k}$ is considered important (denoted as $(u_{i}^{k})^{+}$) if its score $\mathcal{I}(u_{i}^{k})$ exceeds $\delta_{k}$, and unimportant ($(u_{i}^{k})^{-}$) otherwise.
@@
 Only the importance scores of $\hat{\mathcal{U}}_{k-1}$ and $\mathcal{U}_{k}$ are preserved for use between tasks, starting with $\hat{\mathcal{U}}_{1}=\mathcal{U}_{1}$ estimated from $f_{1}$ on $D_{1}$.
-Detailed implementation of entire KlF algorithm is provided in Algorithm 2.
-Algorithm 2 KlF
+Detailed implementation of entire KIF algorithm is provided in Algorithm 2.
+Algorithm 2 KIF
 0:  Dataset $\mathcal{D}_{k}$ for task $k$ = $1,\ldots,K$; initial pre-trained model $f_{0}$; hyperparameters $\alpha_{1},\alpha_{2},\beta,\tau$.
@@
 2:  for task $k$ = $1,\ldots,K$ do
-3:     # knowledge localization.
+3:     # knowledge identification.
 4:     Get $f_{k}$ and calculate $\mathcal{U}_{k}$ by Algorithm (1);
@@
 16:  end for
-V KlF with Memory Replay: KlF-M
-To adapt KlF for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, KlF-M, designed to further improve model performance.
+V KIF with Memory Replay: KIF-M
+To adapt KIF for scenarios where historical data are accessible, we introduce a memory replay-enhanced version, KIF-M, designed to further improve model performance.
 The implementation process is shown in Figure 3.
-Figure 3: Overview of KlF-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of KlF.
-Following the training on task $\mathcal{T}_{k}$ using the knowledge localization and fusion, we obtain the model $\hat{f}_{k}$.
+Figure 3: Overview of KIF-M. By fixing the current task-specific skill unit during the replay phase, we can further enhance the performance of KIF.
+Following the training on task $\mathcal{T}_{k}$ using the knowledge identification and fusion, we obtain the model $\hat{f}_{k}$.
 Before the next task arrives, a replay stage is introduced, where the model $\hat{f}_{k}$ is fine-tuned using historical data stored in the memory buffer $\mathcal{M}_{<k}$.
@@
 O-LoRA [18] 25.89 8.14 -24.59 69.24 10.15 -4.05
-KlF (ours) 26.41 11.78 -18.55 70.71 11.80 -3.67
-KlFLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -3.04
+KIF (ours) 26.41 11.78 -18.55 70.71 11.80 -3.67
+KIFLoRA (ours) 27.43 11.02 -16.91 72.29 12.89 -3.04
 Replay 35.37 2.35 -15.79 70.98 3.28 -15.42
 SAPT [15] 51.54 8.88 -0.57 82.02 9.86 -1.25
-KlF-M (ours) 52.12 11.13 -1.31 84.21 12.32 -1.08
-KlFLoRA-M (ours) 53.01 11.21 -0.81 84.36 11.71 -0.98
+KIF-M (ours) 52.12 11.13 -1.31 84.21 12.32 -1.08
+KIFLoRA-M (ours) 53.01 11.21 -0.81 84.36 11.71 -0.98
 VI-A2 Evaluation Metrics
@@
 VI-A3 Baselines
-We evaluate KlF against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
+We evaluate KIF against the following PEFT-based CL baseline methods: SeqLoRA: sequentially trains the LoRA on the task orders.
 IncLoRA: incremental learning of new LoRA parameters on a sequential series of tasks.
@@
 SAPT [15]: leverages pseudo samples and a shared attention framework to align PEFT block learning and selection.
-Table II compares our KlF with other baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities.
-To clarify the specific localization and fusion techniques used in KlF and its variants: KlF: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for knowledge localization (Eq.
+Table II compares our KIF with other baselines, highlighting three key advantages: data privacy-friendliness, model parameter efficiency, and improved generalization capabilities.
+To clarify the specific identification and fusion techniques used in KIF and its variants: KIF: our foundational framework incorporates matrix-level skill units, utilizes a first-order gradient-based metric for knowledge identification (Eq.
 8), and employs static weighted fusion (Eq.
 15).
-KlFLoRA: an extension of KlF that introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric (Eq.
+KIFLoRA: an extension of KIF that introduces LoRA-tailored skill units, adopts a second-order gradient approximation metric (Eq.
 9), and features adaptive weighted fusion (Eq.
 17).
-KlF-M and KlFLoRA-M: these enhancements of KlF and KlFLoRA integrate memory replay to further boost performance.
-TABLE II: The comparison between KlF and other CL methods.
+KIF-M and KIFLoRA-M: these enhancements of KIF and KIFLoRA integrate memory replay to further boost performance.
+TABLE II: The comparison between KIF and other CL methods.
 Specifically, RF indicates whether the method is rehearsal-free.
@@
 SAPT [15]
-KlF
+KIF
 VI-A4 Implementation Details
 We utilize two distinct language model architectures: the encoder-decoder T5 model [49] and the decoder-only LLaMA model [50].
-In KlF, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq.
+In KIF, the hyperparameters $\alpha_{1}$ and $\alpha_{2}$ in Eq.
 (10) and Eq.
@@
 The overall CL results of various methods using the same T5-large backbone are summarized in Table I.
-Firstly, KlF and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer.
+Firstly, KIF and its variants consistently demonstrate superior CL performance by effectively facilitating knowledge transfer.
 Compared to previous replay-free methods like IncLoRA and O-LoRA, our methods excel in addressing the critical challenges of CF and KT, achieving the highest AP, FWT, and BWT when learning tasks sequentially.
-Furthermore, our enhanced version, KlFLoRA, significantly outperforms KlF.
-Specifically, KlF-M achieves average gains of 1.3% in AP and 0.3% in BWT, underscoring the effectiveness of enhancements to each framework component.
-Notably, even without memory replay, KlF-M surpasses the performance of vanilla repla on the Long Sequence Benchmark, particularly in BWT.
+Furthermore, our enhanced version, KIFLoRA, significantly outperforms KIF.
+Specifically, KIF-M achieves average gains of 1.3% in AP and 0.3% in BWT, underscoring the effectiveness of enhancements to each framework component.
+Notably, even without memory replay, KIF-M surpasses the performance of vanilla repla on the Long Sequence Benchmark, particularly in BWT.
 For the memory-based methods, to ensure a fair comparison, we adopted the same data synthesis-based generative replay approach used in SAPT, ensuring that the amount of data used in memory replay is consistent across methods.
-The results show that both KlF-M and KlFLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework.
-For subsequent experimental analyses and comparisons, we will focus on KlFLoRA, the most effective memory-free version, for further evaluations.
-Secondly, KlFLoRA consistently demonstrates superior performance across various backbones.
+The results show that both KIF-M and KIFLoRA-M surpass the best SAPT models, further proving the versatility and robustness of our framework.
+For subsequent experimental analyses and comparisons, we will focus on KIFLoRA, the most effective memory-free version, for further evaluations.
+Secondly, KIFLoRA consistently demonstrates superior performance across various backbones.
 To further validate our framework’s effectiveness, we conducted experiments using a range of parameter-level backbones.
 As depicted in Figure 4, performance improvements are evident with increasing model size.
-For instance, in T5-XL, KlFLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
+For instance, in T5-XL, KIFLoRA boosts the AP metric from 28.6% to 33.3%, and shows substantial enhancements in FWT and BWT, rising from 11.1% to 16.1% and improving from -20.6% to -14.0%, respectively.
 These results further validate the robust generalization ability of our method.
@@
 (b) Results on the Long Sequence benchmark
-Figure 4: Performance of KlFLoRA across various backbones.
+Figure 4: Performance of KIFLoRA across various backbones.
 Thirdly, our knowledge fusion technique effectively counters catastrophic forgetting.
 To rigorously assess our model’s ability to mitigate forgetting, we evaluated its performance on the initial task after training on subsequent tasks using the Long Sequence Benchmark.
-Figure 5 illustrates that KlFLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task.
+Figure 5 illustrates that KIFLoRA significantly reduces the rate of forgetting, with an average performance decline of only 10% after training on the final task.
 In contrast, vanilla backbones suffer a substantial average performance drop of 22%, highlighting our method’s enhanced capacity to preserve learned knowledge.
@@
 T5-ZS refers to the zero-shot task adaptation methods employed.
-KlFLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
+KIFLoRA exhibited the best performance among all methods tested, a success largely due to its proficient identification and management of task-specific and shared parameters.
 TABLE III: Results on unseen tasks using the T5-Large backbone.
@@
 LFPT5 6.96 35.32 35.00 13.26 21.51 22.41
-KlFLoRA 10.32 31.34 37.13 14.20 47.17 28.03
+KIFLoRA 10.32 31.34 37.13 14.20 47.17 28.03
 VI-D Visualization of Skill Units
@@
   • The distribution of important skill units is task-dependent, illustrating the presence of both task-shared and task-specific parameters.
-This observation substantiates the feasibility of the design motivations behind the KlF framework.
+This observation substantiates the feasibility of the design motivations behind the KIF framework.
   • For classification tasks, such as those in the Long Sequence Benchmark (Figure 6), the skill units in the encoder, particularly the lower layers closer to the model input, play a more pivotal role.
@@
 Figure 7: Visualization of importance scores for LoRA-tailored skill units across different tasks on T5-large within the SuperNI Benchmark.
-TABLE IV: Ablation Study for Importance-Aware Knowledge Localization.
+TABLE IV: Ablation Study for Importance-Aware Knowledge Identification.
 Method AP FWT BWT
@@
 $s\left(\cdot\right)=\left|\nabla_{w_{ij}}\mathcal{L}\right|$ 68.82 10.80 -6.22
-KlFLoRA (ours) 72.29 12.89 -2.04
+KIFLoRA (ours) 72.29 12.89 -2.04
 VI-E Ablation Study
-In this section, we assess the impact of importance-aware knowledge localization, fine-grained knowledge fusion, and memory size on model performance.
+In this section, we assess the impact of importance-aware knowledge identification, fine-grained knowledge fusion, and memory size on model performance.
 Detailed analysis of hyperparameter sensitivity is provided in the Appendix.
-VI-E1 Effect of the Proposed Importance Metric in Knowledge Localization
+VI-E1 Effect of the Proposed Importance Metric in Knowledge Identification
 We explore alternative importance scoring methods to validate the effectiveness of our approach: (i) To evaluate the impact of using moving averages on trajectory gradients, we modify $s(\cdot)$ in Eq.
@@
 Additionally, compared to the first-order gradient approximation, our novel second-order gradient method results in performance improvements of up to 3.4%, 2.1%, and 4.2% across the evaluated metrics.
-These findings underscore the significant role of accurate knowledge localization.
+These findings underscore the significant role of accurate knowledge identification.
 VI-E2 Effect of Adaptive Knowledge Fusion
@@
 EMA 62.76 8.23 -9.80
-KlFLoRA (ours) 72.29 12.89 -3.04
+KIFLoRA (ours) 72.29 12.89 -3.04
 VI-E3 The Effect of Memory Size
-We examine how varying the memory size affects the performance of Replay, KlF-M, and KlFLoRA-M.
+We examine how varying the memory size affects the performance of Replay, KIF-M, and KIFLoRA-M.
 By adjusting the memory size per task $|M|$ to 2%, 5%, 10%, 50%, we summarized the results in Table VI.
@@
 When the memory size is substantially increased, Replay approaches the functionality of multi-task learning, which, while effective, incurs considerable storage and computational costs.
-Conversely, KlF-M and KlFLoRA-M utilize knowledge fusion strategies to effectively maintain parameters that store historical knowledge.
+Conversely, KIF-M and KIFLoRA-M utilize knowledge fusion strategies to effectively maintain parameters that store historical knowledge.
 Consequently, while increasing memory size does boost their performance, the improvements are less dramatic as the memory size grows.
@@
 Replay 71.2 72.4 73.8 76.1
-KlF-M 74.2 75.1 75.7 79.2
-KlFLoRA-M 74.3 75.4 76.2 79.8
+KIF-M 74.2 75.1 75.7 79.2
+KIFLoRA-M 74.3 75.4 76.2 79.8
 VII Conclusion
-In this paper, we introduced the novel Knowledge Localization and Fusion (KlF) framework for language model continual learning.
-KlFLoRA employs importance-aware knowledge localization and fine-grained knowledge fusion to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer.
+In this paper, we introduced the novel Knowledge Identification and Fusion (KIF) framework for language model continual learning.
+KIFLoRA employs importance-aware knowledge identification and fine-grained knowledge fusion to effectively differentiate between task-specific and shared knowledge within each skill unit, mitigating forgetting while enhancing knowledge transfer.
 The framework is highly generalizable and extensible, allowing integration with memory replay methods to boost performance.
-Extensive testing shows that KlFLoRA excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches.
+Extensive testing shows that KIFLoRA excels in preserving past knowledge and performing well in new tasks, surpassing previous state-of-the-art approaches.
 Comprehensive experiments demonstrate demonstrate our framework’s exceptional ability.
@@
 Dynamically selecting these thresholds based on data distribution may yield a more precise classification of task-shared and task-specific parameters, further enhancing performance.
-Additionally, merging the knowledge localization and fusion phases could allow for ongoing parameter consolidation based on their importance during model training.
+Additionally, merging the knowledge identification and fusion phases could allow for ongoing parameter consolidation based on their importance during model training.
 This integrated approach would facilitate more flexible adaptations and better address scenarios in online continual learning.